### v1

# Theory of two-component superfluidity of microcavity polaritons

Abstract
We investigate Bose–Einstein condensation and two-component superfluidity composed of upper (UP) and lower (LP) polaritons confined to a two-dimensional microcavity. By studying a modified Hamiltonian which incorporates interactions between both polariton branches, we derive the collective excitation spectrum and express the sound velocity as a function of detuning and Rabi splitting. Our analysis reveals that the interplay between effective masses and exciton–photon fractions in the UP and LP branches can significantly enhance the superfluid properties compared to a one-component condensate of only lower polaritons. Specifically, we discover that larger Rabi splitting increases the sound velocity and raises the critical temperature for superfluidity, while negative and positive detuning further favors a two-component condensate. Additionally, we demonstrate that these trends are consistent and applicable to various materials, including GaAs and the transition metal dichalcogenides (TMDCs), thereby indicating the potential for high-temperature superfluid phenomena in two-dimensional polariton systems. We hope these findings will provide a deeper understanding of the physics of structures containing multi-component polariton systems and generate experimental realizations of tunable quantum fluids with light.
I Introduction
Bose-Einstein condensation (BEC) and superfluidity in exciton-polariton systems have been emerging as profoundly interesting areas of investigation in condensed matter physics. This is due to the unique properties of polaritons which are quasiparticles arising from the strong coupling between excitons and photons Moskalenko2000 ; Littlewood2007 . These hybrid particles exhibit extremely light effective masses, leading to high critical temperatures for condensation. The critical temperature is relatively high due to the inverse relationship between the 2D thermal de Broglie wavelength and the mass of the quasiparticle. As this wavelength approaches the interparticle spacing, Bose-Einstein condensation (BEC) and superfluidity of exciton-polaritons have been experimentally observed in microcavities Carusotto2013 ; Proukakis2017 ; Peng2022 . Much of the existing research focuses on polaritons in microcavities and under external potentials such as harmonic and anharmonic traps berman2008 .
When a photon is absorbed in a semiconductor an electron is excited from the valence to the conduction band, leaving behind a positively charged vacancy or “hole”. The electron–hole pair can be bound by the Coulomb attraction, thereby forming a quasiparticle which is referred to as an exciton exciton1983 . Under certain densities, these excitons are predicted to exhibit BEC and superfluidity at temperatures significantly higher than those achievable with atomic gases butov2003 . Exciton-polaritons, formed when these excitons strongly couple with photons in an optical microcavity, exhibit both light- and matter-like characteristics deng2010 ; snokeKeeling2017 . This unusual superposition facilitates macroscopic quantum states at relatively high temperatures, making polaritons an attractive means for investigating BEC and superfluidity in two-dimensional (2D) systems Carusotto2013 ; Proukakis2017 ; Peng2022 . Studies have also shown that exciton superfluidity can remain robust even with disorder or magnetic field effects berman2006 . In addition to condensation and superfluidity, polaritons can undergo parametric scattering and exhibit optically parametrized oscillations ciuti2003 .
Previous work on polariton condensation has demonstrated the existence of coherent states in semiconductor microcavities. This has been accompanied by experimental observations in GaAs berman2008 and CdTe at cryogenic temperatures richard2005 . Recent advances in materials science and engineering have expanded these investigations to include monolayer transition metal dichalcogenides (TMDCs) manzeli2017 and organic materials lerario2017 , expanding the realms of polariton condensation.
Two-dimensional van der Waals materials, such as atomically thin transition metal dichalcogenides, possess very distinctive physical properties which make them appealing for a wide range of applications. Monolayers of TMDCs such as MoS2, MoSe2, MoTe2, WS2, WSe2, and WTe2 function as 2D direct-bandgap semiconductors with various applications in electronics and optoelectronics mak2010 . Interest in these monolayers stems from their direct bandgap structure splendiani2010 , the existence of excitonic valley physics, and the feasibility of electrically tunable, strong light–matter interactions xiao2012 . Monolayer TMDCs have already been adopted in field-effect transistors, logic devices, and in both lateral and tunneling optoelectronic designs wang2012 . They exhibit hexagonal lattice structures, with the K and K points in the hexagonal Brillouin zone representing the valleys in the valence and conduction bands Kormanyos2015 . Their excitonic properties have been investigated in numerous experimental and theoretical studies xu2014 . The notable binding energy and extended lifetime of interlayer excitons in van der Waals heterostructures rivera2015 have also motivated substantial research hong2014 . Exciton–polaritons in a TMDC monolayer inside a microcavity were observed experimentally at room temperatures liu2015 , while the superfluidity of exciton-polaritons in such TMDC-based microcavities has been examined in Refs dufferwiel2015 ; kogar2017 . Doping can result in the creation of trions in MoS2, which possess significant binding energies Mak2013 . Furthermore, the ability to create heterostructures by stacking TMDC layers enables the formation of interlayer excitons, which exhibit long lifetimes and robust coherence properties rivera2015 . TMDC-based polaritonic systems embedded in microcavities have shown superfluidity at room temperature, underscoring their potential for future quantum technologies Peng2022 . Besides TMDC double layers, dipolar excitons can also achieve superfluidity in more exotic lattices such as $\alpha$-T3, under appropriate mass and coupling conditions berman2022nanomaterials Exotic phases such as polaritonic or excitonic time crystals have been proposed in TMDC strips subjected to periodic potentials martins2023 , highlighting the breadth of many-body phenomena accessible in 2D materials. Exciton-related phenomena are not limited to conventional semiconductors and TMDCs. Recent studies extend to topological nodal-line semimetals, revealing unique anisotropic optical responses and potential excitonic states in exotic band structures kandel2023 ; kandel2024 .
Gallium Arsenide (or GaAs) is a widely studied III–V semiconductor with a direct bandgap, making it highly suitable for microcavity-based optoelectronic and photonic applications weisbuch1992 . Specifically, GaAs quantum wells embedded within optical microcavities enable the realization of strong exciton-photon coupling and the formation of exciton-polaritons Skol1998 . This strong coupling gives rise to a characteristic Rabi splitting in the excitation spectrum, which underpins the observation of polaritonic effects such as condensation and superfluid-like behavior at comparatively high temperatures Sanvitto2016 . The well-developed epitaxial growth techniques for GaAs allow for precise control over layer thickness, doping profiles, and defect density, ensuring high-quality microcavity structures with minimal disorder steger2015 . These properties, alongside the large oscillator strength and relatively narrow excitonic linewidths in GaAs quantum wells khitrova2006 favor low-threshold polariton lasing zhao2021 and stable Bose–Einstein condensation of polaritons kasprzak2006 . GaAs-based devices serve as a versatile platform for investigating fundamental light–matter interactions, as well as for designing novel photonic components such as polariton lasers, optical switches, and other polariton-based quantum technologies.
Exciton-polaritons (henceforth referred to simply as “polaritons”) form in semiconductor microcavities through the strong coupling of quantum well excitons with photons confined between two Bragg mirrors szymanska2004 . The coupling strength is manifested as Rabi splitting in the excitation spectrum, resulting in two distinct branches: the lower polariton (LP) and the upper polariton (UP). The LP branch has a lower energy and an extremely small effective mass, allowing it to behave as a two-dimensional weakly interacting Bose gas. Experiments have shown that polaritons can exhibit both BEC and superfluidity in microcavity systems snokecoherent2006 . Their tunable interactions, achieved through external fields or cavity-parameter adjustments, render polaritons ideal for studying quantum many-body phenomena and for realizing next-generation photonic devices.
Recent experimental observations Chen2023 have demonstrated Bose–Einstein condensation of upper polaritons in a WS2-based microcavity, showing that under specific detuning and pumping conditions, condensation in the upper polariton branch can become the dominant phase. Additionally, experiments Comaron2025 have highlighted how tunable polariton interactions in driven-dissipative systems enhance spatial coherence and enable Berezinskii–Kosterlitz–Thouless-like crossovers, consistent with interaction-driven superfluid behavior. These developments underscore the importance of incorporating both upper and lower polariton condensates in a unified theoretical description of polariton superfluidity and Bose–Einstein condensation.
In this paper, we direct our attention to a simpler yet equally rich scenario - a two-component condensate of both upper and lower polaritons without the influence of in-plane potentials. We allow scattering between the upper and lower polariton branches, and investigate how this coupling affects the superfluid transition and the resulting collective excitations. By adapting approaches previously applied to two-component exciton superfluidity arafat2024semi , our study provides new insights into multi-component polariton system’s behavior and clarifies how superfluidity can emerge under simpler physical conditions. We further analyze the way in which detuning plays a role in determining the critical temperature and sound velocity, thereby shedding light on the role of multi-component dynamics in polaritonic superfluidity. We find that the Bose gas of polaritons, forming a two-component Bose-Einstein condensate of lower and upper polariton, demonstrate a higher sound velocity and critical temperature rather than just a one-component condensate of lower polaritons. These findings build upon recent advances in polariton condensates and highlight their promise for controlling light-matter interactions in future quantum technologies Sanvitto2016 ; alyatkin2020 ; Bhattacharya2013 . In this work, we show that the effective masses of the upper and lower polariton branches, along with their exciton and photon fractions, vary significantly with detuning, leading to a richer interplay in the two-component condensate than in a one-component condensate. We find that the sound velocity in the two-component condensate is generally higher than in the single-component condensate, and that the critical temperature for superfluidity is also enhanced across all detuning, when other parameters are held constant. Increasing the Rabi splitting consistently raises the sound velocity and therefore the critical temperature, underscoring the benefits of stronger light–matter coupling. These trends hold in both GaAs and various TMDC materials examined, indicating their potential as versatile candidate materials for polariton superfluidity at higher temperatures. Our two-component approach offers enhanced superfluid properties through additional interaction channels and opens new avenues for parameterized engineering in polaritonic devices, potentially enabling high-temperature operation.
The rest of the paper is organized as follows. In Section II, we modify the Hamiltonian for a two-dimensional system of excitons for the Bose gas of polaritons, forming a two-component Bose-Einstein condensate of lower and upper polaritons. We reexpress it in the basis of upper and lower polariton operators, carefully ensuring that we keep the relevant interactions. Details of this process are presented in Appendix B-D. In Section III we derive the spectrum of collective excitations and carefully incorporate the Hopfield coefficients into the spectrum. In Section IV, we build on that and express the sound velocity of the two-component condensate of upper and lower polaritons as a function of these Hopfield coefficients, which allows us to investigate the manner in which the detuning and Rabi splitting affect the sound velocity. In this section we also include our comparisons with the previously investigated one-component picture of solely lower polaritons in GaAs berman2008 . In Section V we develop the subsequent mean-field model of superfluidity and find the critical temperature for superfluidity in the two-component Bose gas of upper and lower polaritons as a function of various parameters. In Section VI, we further our analysis by using various TMDCs each with their own effective masses of excitons. We also compare the two-component condensate and superfluid to a model one-component condensate and superfluid in various TMDCs. In Section VII we present our conclusions and a discussion of our results.
II Effective Hamiltonian of upper and lower polaritons
In this section, we modify the Hamiltonian for our system of excitons and photons to account for both upper and lower polaritons. By analyzing this modified Hamiltonian, we aim to study the critical temperature for Bose-Einstein condensation and the superfluidity of a two-component polariton system, focusing on the role played by scattering between the polariton branches and the collective excitations in the system.
Polaritons are quantum superpositions of excitons and photons. The total Hamiltonian of excitons and microcavity photons is given by $\hat{H}_{total}~{}=~{}\hat{H}_{exc}+\hat{H}_{ph}+\hat{H}_{exc-ph}$, where $\hat{H}_{exc}$ is the excitonic Hamiltonian, $\hat{H}_{ph}$ is the photonic Hamiltonian, and $\hat{H}_{exc-ph}$ is a Hamiltonian of exciton-photon interaction. We can define each of these explicitly.
We begin with the Hamiltonian for 2D excitons in an infinite homogeneous system, given by
$\hat{H}_{\text{exc}}=\sum_{P}\varepsilon_{\text{ex}}(P)\hat{b}_{P}^{\dagger}\hat{b}_{P}+\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}U_{q}\hat{b}_{P+q}^{\dagger}\hat{b}_{P^{\prime}-q}^{\dagger}\hat{b}_{P}\hat{b}_{P^{\prime}}\ ,$ (1)
where $A$ is a normalization area. Here, the first term represents the energy of the excitons, and the second term represents the exciton-exciton interaction. The operator $\hat{b}_{P}$ is the excitonic annihilation operator, and the operator $\hat{b}^{\dagger}_{P}$ is the excitonic creation operator for momentum $P$. The contact potential $U_{q}$ is approximated by Eq. (12). The exciton energy dispersion $\varepsilon_{\text{ex}}(P)$ is given by
$\varepsilon_{\text{ex}}(P)=E_{\text{band}}-E_{\text{binding}}+\varepsilon_{0}(P)\ ,$ (2)
The Hamiltonian for noninteracting microcavity photons is given by
$\hat{H}_{\text{ph}}=\sum_{P}\varepsilon_{\text{ph}}(P)\hat{a}_{P}^{\dagger}\hat{a}_{P}\ \ .$ (3)
Here, the operator $\hat{a}_{P}$ is the photonic annihilation operator, and the operator $\hat{a}^{\dagger}_{P}$ is the photonic creation operator for momentum $P$. For photons confined to a microcavity, the energy dispersion $\varepsilon_{\text{ph}}(P)$ is given by
$\varepsilon_{\text{ph}}(P)=\frac{c}{n}\sqrt{P^{2}+\hbar^{2}\pi^{2}L_{c}^{-2}}\ .$ (4)
where $L_{c}$ is the microcavity length. The exciton-photon interaction term is given by
$\hat{H}_{\text{exc-ph}}=\sum_{P}\hbar\Omega_{R}(\hat{b}_{P}^{\dagger}\hat{a}_{P}+\hat{a}_{P}^{\dagger}\hat{b}_{P}),$ (5)
where $\hbar\Omega_{R}$ is the Rabi splitting, representing the coupling strength between photons and excitons.To fully describe the system, we express the Hamiltonian in the basis of upper and lower polariton operators. We substitute the following expressions from Equation (6) of Ref. hopfield1958 .
$\hat{b}_{P}=X_{P}\hat{l}_{P}-C_{P}\hat{u}_{P},\quad\hat{b}_{P}^{\dagger}=X_{P}\hat{l}_{P}^{\dagger}-C_{P}\hat{u}_{P}^{\dagger}$ (6)
where $\hat{l}_{P}$ is the lower polariton annihilation operator for momentum $P$, $\hat{u}_{P}$ is the upper polariton annihilation operator for momentum $P$, $X_{P}$ is the excitonic fraction of the lower polariton and $C_{P}$ is the photonic fraction of the lower polariton. The photonic operators may be defined in the basis of upper and lower polariton operators as
$\hat{a}_{P}=C_{P}\hat{l}_{P}+X_{P}\hat{u}_{P},\quad\hat{a}_{P}^{\dagger}=C_{P}\hat{l}_{P}^{\dagger}+X_{P}\hat{u}_{P}^{\dagger}.$ (7)
The photon and exciton fractions in the lower polariton branch, denoted as $X_{P}$ and $C_{P}$ respectively, describe the relative contributions of the photonic and excitonic components in the polariton wavefunction. These fractions are given by
$X_{P}=\frac{1}{\sqrt{1+\left(\frac{\hbar\Omega_{R}}{\varepsilon_{\text{LP}}(P)-\varepsilon_{\text{ph}}(P)}\right)^{2}}},$ (8)
$C_{P}=-\frac{\frac{\hbar\Omega_{R}}{\varepsilon_{\text{LP}}(P)-\varepsilon_{\text{ph}}(P)}}{\sqrt{1+\left(\frac{\hbar\Omega_{R}}{\varepsilon_{\text{LP}}(P)-\varepsilon_{\text{ph}}(P)}\right)^{2}}},$ (9)
where $\varepsilon_{\text{LP}}(P)$ is the energy of the lower polariton branch, $\varepsilon_{\text{ph}}(P)$ is the photon dispersion. The polariton energy is given by
$\varepsilon_{\text{LP/UP}}(P)=\frac{\varepsilon_{\text{ph}}(P)+\varepsilon_{\text{exc}}(P)}{2}\mp\frac{1}{2}\sqrt{\left(\varepsilon_{\text{ph}}(P)-\varepsilon_{\text{exc}}(P)\right)^{2}+4|\hbar\Omega_{R}|^{2}}\ ,$ (10)
The linear part of the Hamiltonian, which includes the linear part of $\hat{H}_{\text{exc}}$ and $\hat{H}_{\text{ph}}$, and the interaction between exciton and photon $\hat{H}_{\text{exc-ph}}$ can be diagonalized by applying unitary transformations and has the form
$\hat{H}_{0}=\sum_{\textbf{P}}\varepsilon_{LP}(P)\hat{l}_{\textbf{p}}^{\dagger}\hat{l}_{\textbf{p}}+\sum_{\textbf{P}}\varepsilon_{UP}(P)\hat{u}_{\textbf{p}}^{\dagger}\hat{u}_{\textbf{p}}$ (11)
The next thing we need to consider are interactions. We substitute the expressions for $\hat{b}_{P}$ and $\hat{b}_{P}^{\dagger}$ into the excitonic interaction term,
$\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}U_{q}\hat{b}_{P+q}^{\dagger}\hat{b}_{P^{\prime}-q}^{\dagger}\hat{b}_{P}\hat{b}_{P^{\prime}}$
where, the contact potential $U_{q}$ is approximated by
$U_{q}=\frac{6e^{2}a_{2D}}{\epsilon}$ (12)
and $a_{2D}$ is the effective 2D Bohr radius of excitons,
$a_{2D}=\frac{\hbar^{2}\epsilon}{2\mu_{e-h}e^{2}}$ (13)
This becomes
$\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}U_{q}\left(X_{P+q}\hat{l}_{P+q}^{\dagger}-C_{P+q}\hat{u}_{P+q}^{\dagger}\right)\left(X_{P^{\prime}-q}\hat{l}_{P^{\prime}-q}^{\dagger}-C_{P^{\prime}-q}\hat{u}_{P^{\prime}-q}^{\dagger}\right)\left(X_{P}\hat{l}_{P}-C_{P}\hat{u}_{P}\right)\left(X_{P^{\prime}}\hat{l}_{P^{\prime}}-C_{P^{\prime}}\hat{u}_{P^{\prime}}\right)$ (14)
Expanding this product results in terms involving, pure lower polariton interactions, pure upper polariton interactions, cross-interactions between upper and lower polaritons. We start with the expression
$\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}U_{q}\left(X_{P+q}\hat{l}_{P+q}^{\dagger}-C_{P+q}\hat{u}_{P+q}^{\dagger}\right)\left(X_{P^{\prime}-q}\hat{l}_{P^{\prime}-q}^{\dagger}-C_{P^{\prime}-q}\hat{u}_{P^{\prime}-q}^{\dagger}\right)\left(X_{P}\hat{l}_{P}-C_{P}\hat{u}_{P}\right)\left(X_{P^{\prime}}\hat{l}_{P^{\prime}}-C_{P^{\prime}}\hat{u}_{P^{\prime}}\right)$
The subsequent 16 interaction terms can be combined into a single equation, in a more reasonable order, grouping together terms that are alike. We explicitly sort these terms in Appendix A.
Certain terms in the expanded Hamiltonian represent interactions that are not typical for polaritons due to the following reasons. Polaritons in the upper ($\hat{u}$) and lower ($\hat{l}$) branches generally do not transform into one another without a strong coupling mechanism, especially with the large Rabi splitting. The Rabi splitting represents the energy difference between the upper and lower polariton branches, and when this splitting is significant (on the order of tens of meV), transitions between these branches require a comparable amount of energy, making spontaneous interbranch conversions highly improbable Peng2022 ; Carusotto2013 . As a result, terms where a lower and an upper polariton combine to form two of the same type (either both lower or both upper) are not physically consistent. Processes that would convert one lower and one upper polariton into two lower or two upper polaritons are unlikely to occur without the input of additional energy. This makes certain cross-type conversion terms energetically unfavorable, if even possible, so we discard them from consideration. After removing these non-physical terms, the interaction Hamiltonian simplifies to
$\hat{H}_{\text{int}}=\frac{1}{2A}\sum_{\begin{subarray}{c}\mathbf{P},\mathbf{P}^{\prime},\mathbf{q}\end{subarray}}U_{q}\cdot\frac{1}{4}\Big{[}$ $\hat{l}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}}\hat{l}_{\mathbf{P}^{\prime}}+\hat{u}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}}\hat{u}_{\mathbf{P}^{\prime}}$ (15)
$+\hat{l}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}}\hat{u}_{\mathbf{P}^{\prime}}+\hat{u}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}}\hat{l}_{\mathbf{P}^{\prime}}$
$+\hat{l}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}}\hat{l}_{\mathbf{P}^{\prime}}+\hat{u}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}}\hat{u}_{\mathbf{P}^{\prime}}\Big{]}$
To reiterate, the total Hamiltonian in the upper and lower polariton basis then becomes
$\hat{H}_{\text{total}}$ $=\hat{H}_{0}+\hat{H}_{\text{int}}$ (16)
$=\sum_{\mathbf{P}}\varepsilon_{\text{LP}}(\mathbf{P})\hat{l}_{\mathbf{P}}^{\dagger}\hat{l}_{\mathbf{P}}+\sum_{\mathbf{P}}\varepsilon_{\text{UP}}(\mathbf{P})\hat{u}_{\mathbf{P}}^{\dagger}\hat{u}_{\mathbf{P}}$
$\quad+\frac{1}{2A}\sum_{\begin{subarray}{c}\mathbf{P},\mathbf{P}^{\prime},\mathbf{q}\end{subarray}}U_{q}\cdot\frac{1}{4}\Big{[}\hat{l}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}}\hat{l}_{\mathbf{P}^{\prime}}+\hat{u}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}}\hat{u}_{\mathbf{P}^{\prime}}$
$\qquad+\hat{l}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}}\hat{u}_{\mathbf{P}^{\prime}}+\hat{u}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}}\hat{l}_{\mathbf{P}^{\prime}}$
$\qquad+\hat{l}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}}\hat{l}_{\mathbf{P}^{\prime}}+\hat{u}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}}\hat{u}_{\mathbf{P}^{\prime}}\Big{]}.$
III Spectrum of Collective Excitations
We expect that at $T=0$ K, almost all upper and lower polaritons condense into a BEC of upper and lower polaritons. This allows us to assume the formation of a binary mixture of BECs. We describe this two-component weakly interacting Bose gas of polaritons following the procedure, described in Refs. SuperfluidDipolar ; HighTempSuperOleg ; Pindzola . We then examine this mixture within the Bogoliubov approximation Abrikosov . The collective excitation spectrum for a two-component weakly interacting Bose gas of polaritons is given by
$\Omega_{j,\mathbf{P}}=\sqrt{\frac{\omega_{1,\mathbf{P}}^{2}+\omega_{2,\mathbf{P}}^{2}+(-1)^{j-1}\sqrt{\left(\omega_{1,\mathbf{P}}^{2}-\omega_{2,\mathbf{P}}^{2}\right)^{2}+\left(4G_{12}\right)^{2}\varepsilon_{1,\mathbf{P}}\varepsilon_{2,\mathbf{P}}}}{2}},$ (17)
where $\omega_{j,\mathbf{P}}=\sqrt{\varepsilon_{j,\mathbf{P}}^{2}+2G_{jj}\varepsilon_{j,\mathbf{P}}}$, with $j=1$ for lower polaritons (LP) and $j=2$ for upper polaritons (UP), and their spectra are defined in Eq. (60). Self-interactions are dictated as follows,
$G_{jj}=U_{q}n_{j,0}\ ,$ (18)
[…]
The cross-interaction term is defined as follows,
$G_{12}=U_{q}\sqrt{n_{\text{LP},0}n_{\text{UP},0}}.$ (19)
We expand the spectrum of collective excitations $\Omega_{j,\textbf{P}}$ up to first order with respect to the momentum P and get two modes of the collective excitations $\Omega_{j,\textbf{P}}=c_{j}\textbf{P}$.
$c_{j}=\sqrt{\frac{G_{\text{11}}}{2m_{\text{LP}}}+\frac{G_{\text{22}}}{2m_{\text{UP}}}+(-1)^{j-1}\sqrt{\left(\frac{G_{\text{11}}}{2m_{\text{LP}}}-\frac{G_{\text{22}}}{2m_{\text{UP}}}\right)^{2}+\frac{G_{\text{12}}^{2}}{m_{\text{LP}}m_{\text{UP}}}}}\ ,$ (20)
where $j=1,2$, and $G_{11},G_{22}$ and $G_{12}$ are polariton-polariton interactions defined in Eq. (18) and Eq. (19). The upper and lower polariton masses $m_{\text{UP}}$ and $m_{\text{LP}}$ are defined in Eq. (71) and (70). If we consider the finely tuned and experimentally plausible case where the densities of the polaritons are equal, $n_{\text{LP}}=n_{\text{UP}}=n/2$, we can obtain the sound velocity as
$c_{j}=\sqrt{\frac{U_{q}n}{2}\left(\frac{1}{2m_{\text{LP}}}+\frac{1}{2m_{\text{UP}}}+(-1)^{j-1}\sqrt{\left(\frac{1}{2m_{\text{LP}}}-\frac{1}{2m_{\text{UP}}}\right)^{2}+\frac{1}{m_{\text{LP}}m_{\text{UP}}}}\right)}\ ,$ (21)
It is clear to see that in Eq. (21) there is only one non-zero sound velocity.
$c=\sqrt{\frac{U_{q}n}{2}\left(\frac{1}{m_{\text{UP}}}+\frac{1}{m_{\text{LP}}}\right)}\ .$ (22)
To account for the differences in the concentrations of upper and lower polaritons, we incorporate the Hopfield coefficients $X_{P}$ and $C_{P}$, which represent the excitonic and photonic fractions, respectively. The total Hamiltonian is expressed as,
$\hat{H}_{\text{total}}=\hat{H}_{0}+\hat{H}_{\text{int}}.$ (23)
The non-interacting Hamiltonian is given by
$\hat{H}_{0}=\sum_{\mathbf{P}}\varepsilon_{\text{LP}}(\mathbf{P})\hat{l}_{\mathbf{P}}^{\dagger}\hat{l}_{\mathbf{P}}+\sum_{\mathbf{P}}\varepsilon_{\text{UP}}(\mathbf{P})\hat{u}_{\mathbf{P}}^{\dagger}\hat{u}_{\mathbf{P}},$ (24)
where the energy spectra for lower and upper polaritons are given by Eq. (66). Here, the photon and exciton dispersions are given by Eq. (62)-(63). The Hopfield coefficients $X_{P}$ and $C_{P}$ determine the excitonic and photonic fractions, respectively and we restate them in a convenient form below Griffin1995 ,
$|X_{P}|^{2}=\frac{1}{2}\left(1+\frac{\Delta_{0}}{\sqrt{\Delta_{0}^{2}+4(\hbar\Omega_{R})^{2}}}\right),\quad|C_{P}|^{2}=\frac{1}{2}\left(1-\frac{\Delta_{0}}{\sqrt{\Delta_{0}^{2}+4(\hbar\Omega_{R})^{2}}}\right).$ (25)
We define $\Delta_{0}=\frac{c}{n}\hbar\pi L_{c}^{-1}-\left(E_{\text{band}}-E_{\text{binding}}\right)$ as the detuning. The Hopfield coefficients can be varied widely by varying the detuning and the Rabi splitting as seen below in Fig. 1.
Figure 1: The Hopfield coefficients as a function of the detuning $\Delta_{0}$ defined above for GaAs. We employed the definitions in Eq. (25). We kept the Rabi splitting constant at $\Omega_{R}=$ 15 meV.
It is worth noting that when the detuning is set to 0 meV, the photon and exciton fractions are necessarily fixed at 0.5 each. At this particular detuning level, the Rabi splitting has no effect on the photon and exciton fraction. At any level of non-zero detuning, the fractions are nearly 100% photonic or 100% excitonic. At each of these extremes the detuning is the more prominent parameter than Rabi splitting. However, at a detuning close to, but not quite even, the Rabi splitting has a more prominent effect. This is seen below for GaAs in Fig. 2.
Figure 2: The Hopfield coefficients as a function of the Rabi splitting $\Omega_{R}$ defined above for GaAs. We employed the definitions in Eq. (25). We kept the detuning constant at $\Delta_{0}=$ 10 meV.
When the Rabi splitting is increased, the exciton fraction goes down, and the photon fraction increases. This coupled with the change of the effective masses of the polaritons with the Rabi splitting illustrated in Fig. 20 govern the dynamics of our system, and we will come to see, lead to rich dynamics of the sound velocity and critical temperatures when parametrized by the detuning and the Rabi splitting.
To simplify the notation, we absorb the interaction strength $U_{q}$ into the coefficients $G_{\text{LP}},G_{\text{UP}},$ and $G_{\text{LP-UP}}$. To reflect the asymmetry in polariton populations, we define $n_{\text{LP},0}$ and $n_{\text{UP},0}$ as the total polariton concentrations of lower and upper polaritons, respectively,
$n_{\text{LP},0}=n_{0}|X_{P}|^{2},\quad n_{\text{UP},0}=n_{0}|C_{P}|^{2},$ (26)
where $n_{0}$ is the total polariton concentration. These redefinitions result in the following interaction strengths,
$G_{\text{LP}}=U_{q}n_{0}|X_{P}|^{2},\quad G_{\text{UP}}=U_{q}n_{0}|C_{P}|^{2},\quad G_{\text{LP-UP}}=U_{q}n_{0}|X_{P}||C_{P}|.$ (27)
The interaction Hamiltonian then becomes,
$\hat{H}_{\text{int}}=\frac{1}{2A}\sum_{\begin{subarray}{c}\mathbf{P},\mathbf{P}^{\prime},\mathbf{q}\end{subarray}}\Big{[}G_{\text{LP}}\hat{l}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}}\hat{l}_{\mathbf{P}^{\prime}}+G_{\text{UP}}\hat{u}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}}\hat{u}_{\mathbf{P}^{\prime}}+G_{\text{LP-UP}}\left(\hat{l}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}}\hat{u}_{\mathbf{P}^{\prime}}+\text{.....}\right)\Big{]}.$.. ) ] . (28)
This modifies the interaction terms, leading to observable effects in the collective excitation spectrum and sound velocities, which we derive in the next section.
IV Tunable sound velocity
The collective excitation spectrum for the two-component polariton condensate is given by
$\varepsilon_{j,{P}}=\frac{1}{\sqrt{2}}\left\{\varepsilon_{\text{LP}}({P})^{2}+2G_{\text{LP}}\varepsilon_{\text{LP}}({P})+\varepsilon_{\text{UP}}({P})^{2}+2G_{\text{UP}}\varepsilon_{\text{UP}}({P})+(-1)^{j-1}\sqrt{\beta({P})}\right\}^{1/2}\ ,$ (29)
where
$\beta({P})=\left[\varepsilon_{\text{LP}}({P})^{2}+2G_{\text{LP}}\varepsilon_{\text{LP}}({P})-\varepsilon_{\text{UP}}({P})^{2}-2G_{\text{UP}}\varepsilon_{\text{UP}}({P})\right]^{2}+16G_{\text{LP-UP}}^{2}\varepsilon_{\text{LP}}({P})\varepsilon_{\text{UP}}({P}).$ (30)
We plot this collective excitation spectrum in Fig. 3.
Figure 3: The spectrum of collective excitations in GaAs as a function of momentum. The parameters used in the calculation are $\Omega_{R}=17$ meV, a bare photon of energy 1.6 eV, and a finely adjusted detuning quantity $\Delta_{0}=0$ meV. The collective excitation spectrum is defined in Eq.  (29).
A notable feature worth pointing out about Fig. 3 is the shape of the curves. The collective excitation spectrum for the j = 1 mode is linear in the low momentum regime. This is the excitation which corresponds to the Goldstone mode for superfluidity, and thus the non-zero sound velocity $c_{1}$ in Eq. (39). This mode arises as a result of the spontaneous breaking of the U(1) symmetry of the condensate. Since the condensate prefers a specific phase, small fluctuations in this phase propagate as low-energy excitations, leading to a gapless, sound-like dispersion. In contrast, the j = 2 mode represents a density fluctuation of the condensate.
We can simplify Eq. (29) to give
$\varepsilon_{j,{P}}=\sqrt{\frac{\omega_{\text{LP},{P}}^{2}+\omega_{\text{UP},{P}}^{2}+(-1)^{j-1}\sqrt{\beta({P})}}{2}},$ (31)
where,
$\omega_{\text{LP},{P}}^{2}=\varepsilon_{\text{LP}}({P})^{2}+2G_{\text{LP}}\varepsilon_{\text{LP}}({P}),$ (32)
$\omega_{\text{UP},{P}}^{2}=\varepsilon_{\text{UP}}({P})^{2}+2G_{\text{UP}}\varepsilon_{\text{UP}}({P}),$ (33)
and,
$\beta({P})=\left[\omega_{\text{LP},{P}}^{2}-\omega_{\text{UP},{P}}^{2}\right]^{2}+\left(4G_{\text{LP-UP}}\right)^{2}\varepsilon_{\text{LP}}({P})\varepsilon_{\text{UP}}({P}).$ (34)
Substituting these into $\omega_{\text{LP},{P}}^{2}$ and $\omega_{\text{UP},{P}}^{2}$, we obtain
$\omega_{\text{LP},{P}}^{2}\approx\varepsilon_{\text{LP},0}^{2}+2G_{\text{LP}}\varepsilon_{\text{LP},0}+\frac{P^{2}}{m_{\text{LP}}}\left(\varepsilon_{\text{LP},0}+G_{\text{LP}}\right),$ (35)
$\omega_{\text{UP},{P}}^{2}\approx\varepsilon_{\text{UP},0}^{2}+2G_{\text{UP}}\varepsilon_{\text{UP},0}+\frac{P^{2}}{m_{\text{UP}}}\left(\varepsilon_{\text{UP},0}+G_{\text{UP}}\right).$ (36)
The cross-term $\beta({P})$ contributes
$\beta({P})\approx\text{constant terms}+P^{2}\left(\frac{G_{\text{LP-UP}}^{2}}{m_{\text{LP}}m_{\text{UP}}}\right).$ (37)
Thus, the sound velocity becomes,
$c_{j}=\sqrt{\frac{G_{\text{LP}}}{2m_{\text{LP}}}+\frac{G_{\text{UP}}}{2m_{\text{UP}}}+(-1)^{j-1}\sqrt{\left(\frac{G_{\text{LP}}}{2m_{\text{LP}}}-\frac{G_{\text{UP}}}{2m_{\text{UP}}}\right)^{2}+\frac{G_{\text{LP-UP}}^{2}}{m_{\text{LP}}m_{\text{UP}}}}}.$ (38)
In similar fashion the $j=1$ mode survives while the j = 2 mode leads to $c_{2}=0$, leaving us with the sound velocity of
$c_{1}=c_{s}=\sqrt{\left(\frac{G_{\text{LP}}}{m_{\text{LP}}}+\frac{G_{\text{UP}}}{m_{\text{UP}}}\right)}=\sqrt{U_{q}n_{0}\left(\frac{|X_{P}|^{2}}{m_{\text{LP}}}+\frac{|C_{P}|^{2}}{m_{\text{UP}}}\right)}.$ (39)
We are interested in how the sound velocity varies with the detuning and Rabi splitting. We plot these results below in Fig. 4.
Figure 4: The sound velocity as a function of the detuning in GaAs for different Rabi splittings $\Omega_{R}$. The parameters we kept constant are - total polariton concentration of $n_{0}=10^{8}$ cm-2, and a bare photon energy of 1.6 eV. We employed the definition of the sound velocity $c_{s}$ from Eq. (39).
Fig. 4 illustrates the sound velocity $c_{s}$ as a function of the detuning $\Delta_{0}$ for various values of the Rabi splitting $\Omega_{R}$. The graph reveals that at small detunings, particularly near $\Delta_{0}=0$, the sound velocity is the same regardless of the Rabi splitting. The definition of the effective masses of the polaritons in Eq. (70)-(71) show that at zero detuning, the masses do not vary with the Rabi splitting $\Omega_{R}$. Thus to investigate how the we may vary the sound velocity with the Rabi splitting, we select a detuning in the off-resonance regime. The results are shown below in Fig. 5.
Figure 5: The sound velocity as a function of Rabi splitting in GaAs for chosen detunings. Notably, the plots are symmetric for detunings equidistant from resonance, such that the curves for $+10$ meV and $-10$ meV are identical. The parameters employed are a total polariton concentration of $n_{0}=10^{8}\,\mathrm{cm}^{-2}$ and a bare photon energy of 1.6 eV. The sound velocity $c_{s}$ is determined using the definition in Eq. (39).
We note that in Fig. 5 the sound velocity is increased with increasing Rabi splitting. We may also analyze the differences in the sound velocity between the system with two-components versus one-component Bose-Einstein condensates, using the theoretical framework developed in previous sections. This can enlighten us on the advantages of considering our two-component condensate. The sound velocity in the one-component condensate of lower polariton only is given in Ref. berman2008 and is presented by Eq. (40), i.e.,
$c_{s}=\sqrt{\frac{U_{\text{eff}}n_{0}(r,T)}{M_{\text{eff}}}},$ (40)
In this notation, $U_{\text{eff}}$ is the strength of the effective interaction, $n_{0}(r,T)$ is the total polariton concentration, and $M_{\text{eff}}$ is the effective mass of the lower polariton. The plot in Fig. 6 highlights the distinct behavior of the sound velocity, $c_{s}$, as a function of detuning, $\Delta_{0}$, for both condensates.
Figure 6: (COLOUR ONLINE) The sound velocity as a function of the detuning for our two-component condensate of upper and lower polaritons, and the one-component condensate of lower polaritons only, in GaAs. We use the definition of the sound velocity from Eq. (40) and use the same lower polariton effective mass and the same total concentration of polaritons $n_{0}=10^{8}$ cm-2 in both. We have also included a plot of the lower polariton total polariton concentration as a fraction of the total concentration in both cases. We also used a Rabi splitting of $\Omega_{R}$ = 8 meV.
In Fig. 6, we compare the sound velocity of a polariton Bose gas with a one-component condensate (lower polaritons only) to that of a polariton Bose gas with a two-component condensate, in which the total polariton concentration $n_{0}$ is shared between upper and lower polaritons. Across all detunings, the polariton Bose gas with a two-component condensate exhibits a higher sound velocity. This arises because the two-component condensate benefits from contributions of both polariton branches, whose effective masses vary oppositely with detuning. As shown in Fig. 20(a), while the lower polariton becomes heavier at positive detuning, the upper polariton becomes lighter, and vice versa at negative detuning. This anti-symmetric behavior results in the polariton Bose gas with a two-component condensate displaying a consistently higher sound velocity. The largest difference occurs near zero detuning, where both branches are maximally mixed and their contributions most balanced.
V Superfluidity
In this section, we determine the criterion for which a weakly interacting Bose gas of two components (upper and lower polaritons) can form a superfluid. We exploit the results from the preceding Section IV. Since at low momenta the energy spectrum of quasiparticles of a weakly interacting gas of polaritons is soundlike, this system satisfies the Landau criterion for superfluidity LifPit . The critical velocity for superfluidity is given by $v_{c}=\text{min}(c_{1},c_{2})$ since the quasiparticles are created at velocities exceeding that of sound for the lowest mode of the quasiparticle dispersion.
The density $\rho_{s}(T)$ of the superfluid is defined as $\rho_{s}(T)=\rho-\rho_{n}(T)$, where $\rho=m_{\text{UP}}n_{0}|C_{P}|^{2}+m_{\text{LP}}n_{0}|X_{P}|^{2}$ is the total 2D concentration of the system. Also, $\rho_{n}(T)$ is the density of the normal component. We define the normal component $\rho_{n}(T)$ in the usual way berman2022nanomaterials . We assume that the polariton system moves with velocity u, which means that the superfluid component has velocity u. At finite temperatures $T$, dissipating quasiparticles will emerge in this system. Since their density is small at low temperatures, one may assume that the gas of quasiparticles is an ideal Bose gas. We obtain the density of the superfluid component in the system following the procedure described in Ref. phosphoreneOGR . To calculate the superfluid component density, we begin by defining the mass current J for a Bose gas for quasiparticles in the frame of reference where the superfluid component is at rest by
$\textbf{J}=\int\frac{d^{2}P}{(2\pi\hbar)^{2}}\textbf{P}\left[f[\varepsilon_{1}(\textbf{P})-\textbf{P}\textbf{u}]+f[\varepsilon_{2}(\textbf{P})-\textbf{P}\cdot\textbf{u}]\right]\ ,$ (41)
where $f[\varepsilon_{1}(\textbf{P})]=\{\text{exp}[\varepsilon_{1}(\textbf{P})/(k_{B}T)]-1\}^{-1}$ and $f[\varepsilon_{2}(\textbf{P})]=\{\text{exp}[\varepsilon_{2}(\textbf{P})/(k_{B}T)]-1\}^{-1}$ are the Bose-Einstein distribution functions for the quasiparticles with the angle dependent dispersions $\varepsilon_{1}(\textbf{P})$ and $\varepsilon_{2}(\textbf{P})$ respectively. Expanding the expression under the integral in terms of $\textbf{P}\textbf{u}/k_{B}T$ and restricting ourselves to the first-order term, we obtain
$\textbf{J}=-\frac{\textbf{u}}{2}\int\frac{d^{2}P}{(2\pi\hbar)^{2}}\textbf{P}^{2}\left(\frac{\partial f[\varepsilon_{1}(\textbf{P})]}{\partial\varepsilon_{1}(\textbf{P})}+\frac{\partial f[\varepsilon_{2}(\textbf{P})]}{\partial\varepsilon_{2}(\textbf{P})}\right)\ .$ (42)
Using Eqs.  (41) and(42), along with the definition $\textbf{J}=\rho_{n}\textbf{u}$, we define
$\rho_{n}(T)=-\frac{1}{2}\int\frac{d^{2}P}{(2\pi\hbar)^{2}}\textbf{P}^{2}\left(\frac{\partial f[\varepsilon_{1}(\textbf{P})]}{\partial\varepsilon_{1}(\textbf{P})}+\frac{\partial f[\varepsilon_{2}(\textbf{P})]}{\partial\varepsilon_{2}(\textbf{P})}\right)\ .$ (43)
Exploiting the methods described in Ref. HighTempSuperOleg for the low temperature case, we obtain the mean field critical temperature
$T_{c}=\left[\frac{2\pi\hbar^{2}\rho c_{s}^{4}}{3\zeta(3)sk_{B}^{3}}\right]^{1/3}\ ,$ (44)
where $T_{c}$ is the critical temperature for superfluidity, $\hbar$ is the reduced Planck’s constant, $\rho$ is the total 2D density of the system, accounting for contributions from both polariton branches, $c_{s}$ is the sound velocity defined for the two-component condensate in Eq.  (39) and for the one-component condensate in Eq. (40). Additionally, $\zeta(3)$ is the Riemann zeta function evaluated for a value of $3$, $s$ is the spin degeneracy factor and $k_{B}$ is the Boltzmann constant. For the two component condensate of upper and lower polaritons, this is explicitly written in terms of the Hopfield coefficients and effective masses of the upper and lower polartions as follows
$T_{c}=\left[\frac{2\pi\hbar^{2}\rho}{3\zeta(3)sk_{B}^{3}}\right]^{1/3}\left({U_{q}n_{0}\left(\frac{|X_{P}|^{2}}{m_{\text{LP}}}+\frac{|C_{P}|^{2}}{m_{\text{UP}}}\right)}\right)^{2/3}.$ (45)
For the two component superfluid, we emphasise that $\rho=m_{\text{UP}}n_{0}|C_{P}|^{2}+m_{\text{LP}}n_{0}|X_{P}|^{2}$ is the total 2D density of the system. For comparison we will also be studying how this fares compared to the critical temperature for a one-component superfluid of only lower polaritons
$T_{c}=\left[\frac{2\pi\hbar^{2}\rho}{3\,\zeta(3)\,s\,k_{B}^{3}}\right]^{{1}/{3}}\left(U_{q}\,n_{0}\,\frac{|X_{P}|^{2}}{m_{\mathrm{LP}}}\right)^{2/3}.$ (46)
For this case, it is worth noting that while $n_{0}$ is still the total concentration of polaritons, only the lower polaritons can form a condensate and as such $\rho=m_{\text{LP}}n_{0}|X_{P}|^{2}$ is the correct density to use. We plot the dependance of the critical temperature from Eq. (45) as a function of the detuning in Fig. 7 .
Figure 7: The critical temperature $T_{c}$ as a function of detuning in GaAs for different Rabi splittings. The plots are symmetric with respect to detuning away from resonance, meaning the values for positive and negative detuning (e.g., $\Delta=10\,\text{meV}$ and $\Delta=-10\,\text{meV}$) are identical. The parameters used are a total polariton concentration of $n_{0}=10^{10}\,\text{cm}^{-2}$ and a bare photon energy of  1.6 eV. The critical temperature was calculated using the definition from Eq. (45).
The plot highlights the symmetric nature of $T_{c}$ with respect to $\Delta_{0}=0$, meaning that the critical temperature profiles for positive and negative detuning are identical. The maximum $T_{c}$ occurs at resonance ($\Delta_{0}=0$), where the polaritonic system is most balanced between its excitonic and photonic components. The critical temperature is fairly stable around zero detuning, and exhibits a plateau around that region.
Detuning away from resonance generally leads to a lower critical temperatures. This trend reflects the increased dominance of either the excitonic or photonic fraction in the polariton composition, which causes a lower $T_{c}$. However, the Rabi splitting allows us to change the critical temperature. A larger Rabi splitting leads to larger critical temperature for all materials, across all non-zero detuning levels.
Figure 8: (COLOUR ONLINE) Critical temperature ($T_{c}$) as a function of detuning ($\Delta_{0}$) for one-component and two-component polariton superfluids in GaAs. The one-component superfluid (dashed black line) has a smaller critical temperature than the two-component superfluid for negative detuning, but larger critical temperature past a certain positive detuning. The two-component superfluid (solid red line) shows a pronounced peak around zero detuning, followed by symmetric decreases for positive and negative detuning values. The parameters used were a Rabi splitting $\Omega_{R}=15$ meV, total polariton concentration $n_{0}=10^{10}$ cm-2, and a bare photon energy of 1.6 eV. The critical temperatures were calculated using the definitions from Eq. (44) and Eq. (45).
The two-component superfluid exhibits a peak around zero detuning due to balanced contributions from both lower and upper polaritons. Away from resonance, the detuning-dependent Hopfield coefficients redistribute the weights between the polariton branches, leading to symmetric reductions in the critical temperature for the two-component superfluid. These reductions highlight the diminished contributions from interbranch interactions in the two-component superfluid as the system moves further into excitonic or photonic regimes. It is worth noting that for all $\Delta_{0}$ the one-component superfluid has a lower critical temperature than the two-component superfluid. We are also interested in how the critical temperature for our newly developed two-component superfluid compares with the one-component as a function of Rabi splitting in Fig. 9. It is worth commenting about why the two-component superfluid demonstrates a critical temperature for superfluidity that is symmetric around the detuning $\Delta_{0}=0$, while the one-component superfluid demonstrates a peak away from $\Delta_{0}=0$. From Fig. 1, Fig. 6 and Fig. 20 (a), we may arrive at the answer. From Fig. 6 and Eq. (40), we see that the sound velocity is an even function with respect to the detuning. However both the excitonic fraction from Fig. 1 and the effective mass of the lower polariton from Fig. 20 (a) are both not even functions of detuning, and unlike the two-component superfluid, is not cancelled out by symmetric contributions from the upper polariton branch.
Figure 9: Plot of the critical temperature, $T_{c}$ for the one-component and two-component Bose-Einstein condensate, as a function of the Rabi splitting $\Omega_{R}$ in GaAs. The two plots differ in the following way - plot (a) has a negative detuning of $\Delta_{0}=~{}-10$ meV, and plot (b) has a positive detuning of $\Delta_{0}=10$ meV. The parameters used were a total polariton concentration $n_{0}=10^{10}$ cm-2, and a bare photon energy of 1.6 eV. The critical temperatures were calculated using the definitions from Eq. (44) and Eq. (45).
a) Refer to caption b) Refer to caption
Based on our results in Fig. 9, it is worth restating that for all detuning, the one-component superfluid has a lower critical temperature than the two-component superfluid due to the increased photonic nature of the system, and thus favoring the production of upper polaritons over lower polaritons. We would also like to investigate how the sound velocity and critical temperature vary with the condensate densities, $n_{0}$. In Fig. 10 we plot these results.
Figure 10: Plot of the sound velocity, $c_{s}$ for the one-component and two-polariton condensates, as a function of the total polariton concentration $n_{0}$ in GaAs. The plot represents a positive detuning of $\Delta_{0}=10$ meV. Other parameters used were a Rabi splitting $\Omega_{R}=10$ meV, and a bare photon energy of 1.6 eV.
From Fig. 10, we observe that the sound velocity increases with total polariton concentration. The one-component condensate occupies only the lower polariton branch, with interactions scaled by the square of the excitonic fraction $|X_{P}|^{2}$, and dispersion governed by the effective mass $m_{\text{LP}}$. In contrast, the two-component condensate distributes the same total concentration between both upper and lower polariton branches, with contributions weighted by the Hopfield coefficients. Across the entire concentration range, the two-component condensate exhibits a consistently higher sound velocity, due to the combined effects of both branches and their differing masses and interaction strengths. This emphasizes how including both polariton branches—rather than only one—can enhance collective sound modes in the system. Furthermore, we may investigate how the critical temperature varies with the total polariton concentration. The results are presented in Fig. 11.
Figure 11: Plot of the critical temperature, $T_{c}$ for the one-component and two-polariton superfluids, as a function of the total polariton concentration $n_{0}$ in GaAs. The two plots differ in the following way - plot (a) has a negative detuning of $\Delta_{0}~{}=~{}-10$ meV, and plot (b) has a positive detuning of $\Delta_{0}=10$ meV. The parameters used were a Rabi splitting $\Omega_{R}=10$ meV, and a bare photon energy of 1.6 eV. The critical temperatures were calculated using the definitions from Eq. (44) and Eq. (45).
a) Refer to caption b) Refer to caption
For both fixed positive and negative detuning, and at fixed Rabi splitting, the one-component superfluid consistently exhibits a lower critical temperature for superfluidity compared to the two-component superfluid. At higher total polariton concentrations, the difference in critical temperature between the one- and two-component polariton superfluids becomes more pronounced, with the two-component superfluid exhibiting a significantly higher $T_{c}$ across the full range.
VI Results for transition metal dichalcogenides
We are also interested in seeing how this analysis holds for transition metal dichalcogenides. In this section, we apply our theoretical framework to TMDCs and plot the results in Fig. 12.
Figure 12: The sound velocity $c_{s}$ as a function of detuning $\Delta_{0}$ for various TMDC materials: MoS2, MoSe2, MoTe2, WS2, and WSe2. The sound velocity is symmetric about $\Delta_{0}=0$, where the detuning corresponds to resonance. TMDCs with lighter exciton effective masses, such as WS2 and WSe2, exhibit higher sound velocities, while heavier effective masses, such as MoTe2, lead to lower sound velocities. The parameters used include a total polariton concentration $n_{0}=10^{12}\,\mathrm{cm}^{-2}$, a bare photon energy of 1.6 eV, and a Rabi splitting $\Omega_{R}=5\,\mathrm{meV}$. The sound velocity is calculated using Eq. (39).
Each curve exhibits a distinctive peak at zero detuning, reflecting the crossover between predominantly excitonic and photonic character in the polariton branches. The magnitude and width of these peaks differ among the materials. Far from zero detuning, the sound velocity approaches a steady value for each TMDC. We may want to examine how the sound velocity, $c_{s}$, changes for chosen detuning, for a range of Rabi splittings. The Rabi splittings can vary widely among TMDCs liu2015 ; dufferwiel2015 ; flatten2016 ; wang2018 and so that would be of immense interest to us. We present the results in Fig.  13.
Figure 13: The sound velocity $c_{s}$ as a function of the Rabi splitting $\Omega_{R}$ for different TMDC monolayers (MoS2, MoSe2, MoTe2, WS2, and WSe2) at a fixed detuning of $\Delta_{0}=10$ meV. The plots demonstrate that the sound velocity increases with increasing Rabi splitting for all TMDCs. The variation between the TMDCs reflects differences in the exciton effective masses, which directly influence the polariton effective masses and interaction strengths. We employ the definition of the sound velocity from Eq. (39).
The results presented in Fig. 13 show the sound velocity $c_{s}$ as a function of the Rabi splitting $\Omega_{R}$ for various TMDC monolayers, specifically MoS2, MoSe2, MoTe2, WS2, and WSe2, for chosen detuning parameter $\Delta_{0}=10.0$ meV. The trend indicates that the sound velocity is increased with increasing Rabi splitting for all these materials. This behavior is due to the fact that the fact that larger Rabi splitting enhances the coupling between the exciton and photon components, leading to changes in the effective polariton masses and interaction strengths. This is in agreement with our results for GaAs in Fig. 5. We may also analyze the critical temperature for superfludity for these TMDCs.
Figure 14: Critical temperature $T_{c}$ as a function of detuning $\Delta_{0}$ for various TMDCs at a fixed Rabi splitting of $\Omega_{R}=8$ meV. The plot illustrates the symmetric dependence of $T_{c}$ on detuning, with maximum occurring at resonance ($\Delta_{0}=0$). The distinct curves for each TMDC highlight the impact of differing exciton effective masses, with lighter effective masses (e.g., WS2 and WSe2) resulting in higher critical temperatures. The parameters used include a total polariton concentration of $n_{0}=10^{10}$ cm-2 and a bare photon energy of 1.6 eV. The critical temperatures were calculated using the definition from Eq. (45).
The critical temperature ($T_{c}$) as a function of detuning for different TMDCs is presented in Fig. 15. At zero detuning ($\Delta_{0}=0$), all TMDCs exhibit their maximum $T_{c}$. As the detuning shifts away from zero, the critical temperature decreases symmetrically for both positive and negative detuning values. Notably, TMDCs with lighter exciton effective masses, such as WS2 and WSe2, achieve higher critical temperatures compared to those with heavier exciton masses, such as MoTe2. This highlights the significant influence of material-specific properties in determining the critical temperature, emphasizing the tunability of polaritonic systems for different TMDCs. More importantly, we would like to vary the Rabi splitting for chosen detuning.
Figure 15: Critical temperature $T_{c}$ as a function of Rabi splitting $\Omega_{R}$ for various TMDCs at a fixed detunings of $\Delta_{0}=10$ meV. The plot showcases the increases in critical temperature for fixed detuning of 10 meV. The distinct curves for each TMDC highlight the impact of differing exciton effective masses, with lighter effective masses (e.g., WS2 and WSe2) resulting in higher critical temperatures. The parameters used include a total polariton concentration of $n_{0}=10^{10}$ cm-2 and a bare photon energy of 1.6 eV. The critical temperatures were calculated using the definitions from Eq. (44) and Eq. (45).
We are also interested in examining how the one-component Bose-Einstein condensate of only lower polaritons, compares with the two-component Bose-Einstein condensate of lower and upper polaritons, when we vary the detuning and the Rabi splitting. The results for the critical temperature versus detuning for this scenario are plotted for four TMDCs of our choice in Fig. 16.
Figure 16: Plots of the critical temperature $T_{c}$ for different TMDCs as a function of the detuning, $\Delta_{0}$ for the two-component condensate and the one-component condensate. Each subfigure corresponds to a specific TMDC - (a) MoS2, (b) MoSe2, (c) WS2, and (d) WSe2. The parameters used are a fixed Rabi splitting $\Omega_{R}=10$ meV, a total polariton concentration of $n_{0}=10^{10}\,\mathrm{cm}^{-2}$ and a bare photon energy of 1.6 eV. The critical temperatures were calculated using the definitions from Eq. (44) and Eq. (45).
a) [Uncaptioned image] b) [Uncaptioned image]
c) [Uncaptioned image] d) [Uncaptioned image]
We are also interested in how the critical temperature varies with the Rabi splitting in these TMDCs and we plot the results below. From Fig. 16, we observe that the critical temperature as a function of detuning is consistently higher for all selected TMDCs in the two-component superfluid compared to the one-component superfluid consisting solely of lower polaritons, at fixed Rabi splitting and total polariton concentration. At negative detuning, the two-component superfluid shows a significantly higher critical temperature. Although at positive detuning the one-component superfluid eventually surpasses its negative-detuning counterpart in critical temperature, it remains lower than that of the two-component superfluid across the entire range. These results are consistent with our earlier findings for GaAs in Fig. 8.
Figure 17: Plot of the critical temperature, $T_{c}$ for the one-component and two-component Bose-Einstein condensate, as a function of the Rabi splitting $\Omega_{R}$ in WS2. The two plots differ in the following way - plot (a) has a negative detuning of $\Delta_{0}=-10$ meV, and plot (b) has a positive detuning of $\Delta_{0}=10$ meV. The parameters used were a total polariton concentration $n_{0}=10^{10}$ cm-2, and a bare photon energy of 1.6 eV. The critical temperatures were calculated using the definitions from Eq. (44) and Eq. (45).
a) Refer to caption b) Refer to caption
The results in Fig. 17 are consistent with those obtained for GaAs in Fig. 9. The critical temperature $T_{c}$ for superfluidity is higher in the two-component superfluid of upper and lower polaritons for all fixed negative detunings and varying Rabi splittings $\Omega_{R}$. At fixed positive detuning, the one-component superfluid of lower polaritons yields a higher $T_{c}$ at low Rabi splittings, but as $\Omega_{R}$ increases, the two-component superfluid eventually exhibits a higher critical temperature. In all cases, the total polariton concentration is held fixed. We also investigate how varying the total polariton concentration influences the critical temperature for superfluidity.
Figure 18: Plot of the critical temperature, $T_{c}$ for the one-component and two-polariton condensates, as a function of the total polariton concentration $n_{0}$ in a choice TMDC - MoS2. The two plots differ in the following way - plot (a) has a negative detuning of $\Delta_{0}=-10$ meV, and plot (b) has a positive detuning of $\Delta_{0}=10$ meV. The parameters used were a Rabi splitting $\Omega_{R}=10$ meV, and a bare photon energy of 1.6 eV. The critical temperatures were calculated using the definitions from Eq. (44) and Eq. (45).
a) Refer to caption b) Refer to caption
From Fig. 18, we observe that the two-component superfluid consistently exhibits a higher critical temperature compared to the one-component superfluid across all concentrations and detuning values. This trend mirrors the behavior seen in GaAs, as shown in Fig. 11, where specific choices of detuning and Rabi splitting can lead to a one-component superfluid surpassing the two-component superfluid in critical temperature. While the one-component superfluid may achieve a higher critical temperature under carefully tuned conditions, the two-component superfluid generally maintains a greater critical temperature for superfluidity, particularly at negative detuning and for small positive detuning, in both TMDCs and GaAs.
VII Conclusion and discussion
We have investigated a two-component condensate and superfluid of upper and lower polaritons in semiconductor microcavities and compared our results to those for a one-component condensate and superfluid of lower polaritons only. By retaining the interplay of exciton and photon fractions in both branches, we demonstrated that the effective masses, detuning, and Rabi splitting collectively govern the superfluid behavior of the condensate. Specifically, we have found that the two-component condensate generally exhibits a higher sound velocity and enhanced critical temperature for superfluidity across all detuning. While the one-component superfluid can exhibit a peak in the critical temperature at certain positive detunings, it consistently yields a lower $T_{c}$ compared to the two-component superfluid across the parameter range studied. Increasing the Rabi splitting consistently raises both the sound velocity and critical temperature, underscoring how stronger light–matter coupling can increase the critical temperature for superfluidity.
Our numerical calculations for GaAs quantum wells and various transition metal dichalcogenides (TMDCs) confirm the generality of these trends, indicating the potential of two-component polariton condensates for high-temperature operation across multiple materials. These findings highlight the versatility of multi-branch polariton systems in designing future photonic devices and quantum technologies. By varying the detuning and Rabi splitting, polaritonic devices could utilize the improved stability and tunability offered by two-component condensates. Further theoretical and experimental investigations of spin degrees of freedom, nonlinear interactions, and topological effects may reveal yet richer collective phenomena in exciton–polaritons, paving the way for more advanced applications in coherent light–matter technologies.
Acknowledgements
A. Nafis Arafat and Godfrey Gumbs gratefully acknowledge funding from the U.S. National Aeronautics and Space Administration (NASA) via the NASA-Hunter College Center for Advanced Energy Storage for Space under cooperative agreement 80NSSC24M0177. Oleg Berman acknowledges the support from PSC CUNY Award No. 66382-00 54. This research benefited from Physics Frontier Center for Living Systems funded by the National Science Foundation (PHY- 2317138). The authors are grateful to David W. Snoke for valuable discussions.
Appendix A Systematic interaction reduction
The Hamiltonian interaction terms as we left them were somewhat cumbersome. We outline in this section how we systematically reduce them to obtain the simplified form in Eq. (16). Let us begin with the 16 interaction terms
$\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}U_{q}\Bigg{[}$ $X_{P+q}X_{P^{\prime}-q}X_{P}X_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}-X_{P+q}X_{P^{\prime}-q}C_{P}X_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$ (47)
$-X_{P+q}X_{P^{\prime}-q}X_{P}C_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}+X_{P+q}X_{P^{\prime}-q}C_{P}C_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}$
$-X_{P+q}C_{P^{\prime}-q}X_{P}X_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}+X_{P+q}C_{P^{\prime}-q}C_{P}X_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$-X_{P+q}C_{P^{\prime}-q}X_{P}C_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}+X_{P+q}C_{P^{\prime}-q}C_{P}C_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}$
$-C_{P+q}X_{P^{\prime}-q}X_{P}X_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}+C_{P+q}X_{P^{\prime}-q}C_{P}X_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$-C_{P+q}X_{P^{\prime}-q}X_{P}C_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}+C_{P+q}X_{P^{\prime}-q}C_{P}C_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}$
$+C_{P+q}C_{P^{\prime}-q}X_{P}X_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}-C_{P+q}C_{P^{\prime}-q}C_{P}X_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$-C_{P+q}C_{P^{\prime}-q}X_{P}C_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}+C_{P+q}C_{P^{\prime}-q}C_{P}C_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}\Bigg{]}.$
There are only two straightforward self-interaction terms for the upper and lower polaritons. These are
$X_{P+q}X_{P^{\prime}-q}X_{P}X_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}$ (48)
$C_{P+q}C_{P^{\prime}-q}C_{P}C_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}$ (49)
To simplify and rearrange the 16 interaction terms, we will utilize commutation relations for the bosonic creation and annihilation operators Abrikosov , as well as symmetry properties of the terms. The commutation relations reflect the nature of bosons, which can occupy the same quantum state without any restriction. The basic commutation relations for bosonic operators are,
$[\hat{l}_{P},\hat{l}_{P^{\prime}}^{\dagger}]=\delta_{P,P^{\prime}}$ (50)
$[\hat{u}_{P},\hat{u}_{P^{\prime}}^{\dagger}]=\delta_{P,P^{\prime}}$ (51)
$[\hat{l}_{P},\hat{l}_{P^{\prime}}]=0,\quad[\hat{u}_{P},\hat{u}_{P^{\prime}}]=0$ (52)
where $\hat{l}_{P}$ and $\hat{u}_{P}$ are the annihilation operators for different polariton states. Additionally, since these operators correspond to different modes, their cross-commutators will also be zero:
$[\hat{l}_{P},\hat{u}_{P^{\prime}}]=[\hat{l}_{P},\hat{u}_{P^{\prime}}^{\dagger}]=0$ (53)
Let us set aside the self interaction terms Eq. (48)-(49) for now. We categorize the 14 remaining cross-interaction terms into two main groups based on the operator structures and then apply normal ordering where possible. Terms Involving $\hat{l}^{\dagger}\hat{l}\hat{u}^{\dagger}\hat{u}$ or $\hat{u}^{\dagger}\hat{u}\hat{l}^{\dagger}\hat{l}$ are interesting to us for considering physical scattering processes.
The terms that fit this structure are
$X_{P+q}C_{P^{\prime}-q}X_{P}C_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P^{\prime}}$
$+C_{P+q}X_{P^{\prime}-q}C_{P}X_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P^{\prime}}$
These two terms involve interactions where two polaritons of different types (upper and lower) scatter with respect to each other. The terms that are remaining involve various other combinations of processes. These remaining 12 cross-interaction terms we categorize into two main groups based on the nature of the exchange processes: direct exchange terms and mixed exchange terms. The direct exchange terms represent direct interactions where the upper polariton exchanges with the lower polariton.
$X_{P+q}C_{P^{\prime}-q}C_{P}X_{P^{\prime}}\,\hat{l}_{P+q}^{\dagger}\,\hat{u}_{P^{\prime}-q}^{\dagger}\,\hat{u}_{P}\,\hat{l}_{P^{\prime}}$
$\quad+C_{P+q}X_{P^{\prime}-q}X_{P}C_{P^{\prime}}\,\hat{u}_{P+q}^{\dagger}\,\hat{l}_{P^{\prime}-q}^{\dagger}\,\hat{l}_{P}\,\hat{u}_{P^{\prime}}$ (54)
These two terms describe simple polariton exchange interactions. The following mixed terms involve more complex combinations of operators that we would like to sort through
$X_{P+q}C_{P^{\prime}-q}X_{P}X_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}$
$+X_{P+q}C_{P^{\prime}-q}C_{P}C_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}$
$+C_{P+q}X_{P^{\prime}-q}C_{P}C_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}$
$+C_{P+q}X_{P^{\prime}-q}X_{P}X_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}$
$+X_{P+q}X_{P^{\prime}-q}C_{P}X_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$+X_{P+q}X_{P^{\prime}-q}X_{P}C_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}$
$+C_{P+q}C_{P^{\prime}-q}X_{P}C_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}$
$+C_{P+q}C_{P^{\prime}-q}X_{P}X_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}$
$+X_{P+q}C_{P^{\prime}-q}C_{P}X_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$+C_{P+q}X_{P^{\prime}-q}X_{P}C_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}$
$+X_{P+q}C_{P^{\prime}-q}C_{P}X_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$+C_{P+q}X_{P^{\prime}-q}X_{P}C_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}$ (55)
The terms in Eq. (55) involve three creation/annihilation operators of one type and one of the other. The interactions still involve four polaritons, but the arrangement is asymmetric with respect to ratio of upper and lower polaritons, and extremely unlikely to occur given the Rabi splitting.
Appendix B Hopfield coefficients simplification
The photon and exciton fractions in the lower polariton branch, denoted as $X_{P}$ and $C_{P}$ respectively, describe the relative contributions of the photonic and excitonic components in the polariton wavefunction. These fractions are given by
$X_{P}=\frac{1}{\sqrt{1+\left(\frac{\hbar\Omega_{R}}{\varepsilon_{\text{LP}}(P)-\varepsilon_{\text{ph}}(P)}\right)^{2}}},$ (56)
$C_{P}=-\frac{\frac{\hbar\Omega_{R}}{\varepsilon_{\text{LP}}(P)-\varepsilon_{\text{ph}}(P)}}{\sqrt{1+\left(\frac{\hbar\Omega_{R}}{\varepsilon_{\text{LP}}(P)-\varepsilon_{\text{ph}}(P)}\right)^{2}}},$ (57)
where $\varepsilon_{\text{LP}}(P)$ is the energy of the lower polariton branch, $\varepsilon_{\text{ph}}(P)$ is the photon dispersion, $\hbar\Omega_{R}$ is the Rabi splitting, representing the coupling strength between photons and excitons. Under certain conditions, it is possible to approximate $X_{P}\approx\frac{1}{\sqrt{2}}$, indicating an equal mixture of photonic and excitonic components in the lower polariton state. The approximation $X_{P}\approx\frac{1}{\sqrt{2}}$ holds when the term inside the square root in the definition of $X_{P}$ is equal to 1. Specifically, this occurs when,
$\varepsilon_{\text{LP}}(P)-\varepsilon_{\text{ph}}(P)=\hbar\Omega_{R}.$ (58)
This relationship implies that the energy difference between the lower polariton and the photon at momentum $P$ matches the Rabi coupling strength $\hbar\Omega_{R}$. To take advantage of this approximation the photon and exciton energies should be close to resonance at $P=0$, meaning that the detuning $\Delta_{0}$ is small and this is experimentally adjustable dufferwiel2015 . This allows the lower polariton energy $\varepsilon_{\text{LP}}(0)$ to be approximately halfway between the photon and exciton energies at $P=0$. The Rabi splitting $\Omega_{R}$ should be large enough relative to any small detuning between $\varepsilon_{\text{ph}}(P)$ and $\varepsilon_{\text{exc}}(P)$. This ensures significant mixing of photon and exciton states, leading to equal contributions in the lower polariton. Under these conditions, both $X_{P}$ and $C_{P}$ approach $\frac{1}{\sqrt{2}}$, indicating that the lower polariton state is an equal superposition of photon and exciton components. This simplification is useful because it allows us to treat the lower polariton as an idealized mixture, facilitating further calculations without loss of significant accuracy in this regime.
Using the approximation $X_{P}\approx\frac{1}{\sqrt{2}}$ is justified in many physical scenarios where the detuning is minimal, and the coupling is strong, as it reflects the experimentally relevant case of a balanced polariton. By assuming equal fractions, we simplify our analysis and capture the core physics of the coupled photon-exciton system in the microcavity. This approach is commonly used in polariton studies when the lower polariton branch is at resonance or near-resonance, making it both theoretically and experimentally meaningful to approximate $X_{P}\approx\frac{1}{\sqrt{2}}$.
Appendix C Expanded Polariton Spectrum and Effective Masses
Polaritons, which result from the strong coupling between excitons and photons, have an energy spectrum that reflects the combined properties of both components. To understand the behavior of polaritons, it is essential to compute their energy spectrum $\varepsilon_{\text{LP/UP}}(P)$, where LP and UP refer to the lower and upper polariton branches, respectively. This spectrum depends on both the photon dispersion $\varepsilon_{\text{ph}}(P)$ and the excitonic dispersion $\varepsilon_{\text{exc}}(P)$.
By analyzing the polariton spectrum in momentum space, we can extract the effective masses of the polaritons, which determine how they behave in real space under external potentials or forces. The effective mass $m_{\text{eff}}$ for polaritons can be obtained by taking the second derivative of the energy spectrum $\varepsilon_{\text{LP/UP}}(P)$ with respect to momentum $P$:
$m_{\text{eff}}=\left(\frac{\partial^{2}\varepsilon_{\text{LP/UP}}(P)}{\partial P^{2}}\right)^{-1}$ (59)
Here, $\varepsilon_{\text{LP/UP}}(P)$ is the energy of the lower or upper polariton, and $P$ is the momentum of the polariton. Once the effective mass is known, we can model the polaritons in real space. The polariton energy is given by
$\varepsilon_{\text{LP/UP}}(P)=\frac{\varepsilon_{\text{ph}}(P)+\varepsilon_{\text{exc}}(P)}{2}\mp\frac{1}{2}\sqrt{\left(\varepsilon_{\text{ph}}(P)-\varepsilon_{\text{exc}}(P)\right)^{2}+4|\hbar\Omega_{R}^{2}|}\ ,$ (60)
Where, $\varepsilon_{\text{LP}}(P)$ and $\varepsilon_{\text{UP}}(P)$ are the lower and upper polariton energy dipsersions, respectively, $\varepsilon_{\text{ph}}(P)$ is the photon energy dispersion, $\varepsilon_{\text{exc}}(P)$ is the exciton energy dispersion, $\Omega_{R}$ is the Rabi splitting, describing the coupling strength between the exciton and photon.
For photons confined to a microcavity, the energy dispersion is given by
$\varepsilon_{\text{ph}}(P)=\frac{c}{n}\sqrt{P^{2}+\hbar^{2}\pi^{2}L_{c}^{-2}}\ .$ (61)
Where, $c$ is the speed of light in vacuum, $n$ is the refractive index of the material inside the cavity, $P$ is the in-plane momentum of the photon, $\hbar$ is the reduced Planck’s constant, $L_{c}$ is the length of the cavity in the direction perpendicular to the plane, $\hbar\pi L_{c}^{-1}$ represents the quantized photon momentum due to the confinement.
The exciton energy dispersion is given by
$\varepsilon_{\text{ex}}(P)=E_{\text{band}}-E_{\text{binding}}+\varepsilon_{0}(P)\ .$ (62)
Where, $E_{\text{band}}$ is the band gap energy, $E_{\text{binding}}$ is the exciton binding energy, $\varepsilon_{0}(P)$ is the kinetic energy of the exciton, given by
$\varepsilon_{0}(P)=\frac{P^{2}}{2M_{\text{exc}}}$ (63)
$M_{\text{exc}}$ is the effective mass of the exciton ($M_{\text{exc}}=m_{e}+m_{h}$), and $P$ is the in-plane momentum of the exciton. The photon energy can be approximated to first-order as
$\varepsilon_{\text{ph}}(P)\approx\frac{c}{n}\left(\hbar\pi L_{c}^{-1}+\frac{P^{2}}{2\hbar\pi L_{c}^{-1}}\right)$ (64)
Now we let $\Delta_{0}=\frac{c}{n}\hbar\pi L_{c}^{-1}-\left(E_{\text{band}}-E_{\text{binding}}\right)$.
The expanded square root term is
$\sqrt{\Delta_{0}^{2}+4\Omega_{R}^{2}+2\Delta_{0}\frac{P^{2}}{2}\left(\frac{c}{n\hbar\pi L_{c}^{-1}}-\frac{1}{M_{\text{exc}}}\right)}\approx\sqrt{\Delta_{0}^{2}+4\Omega_{R}^{2}}+\frac{\Delta_{0}}{\sqrt{\Delta_{0}^{2}+4\Omega_{R}^{2}}}\frac{P^{2}}{2}\left(\frac{c}{n\hbar\pi L_{c}^{-1}}-\frac{1}{M_{\text{exc}}}\right)$ (65)
The Taylor-expanded polariton energies are
$\varepsilon_{\text{LP/UP}}(P)\approx$ $\;\frac{1}{2}\left(\frac{c}{n}\hbar\pi L_{c}^{-1}+E_{\text{band}}-E_{\text{binding}}\right)\mp\frac{1}{2}\sqrt{\Delta_{0}^{2}+4\Omega_{R}^{2}}$ (66)
$+\frac{P^{2}}{4}\left[\left(\frac{c}{n\hbar\pi L_{c}^{-1}}+\frac{1}{M_{\text{exc}}}\right)\mp\frac{\Delta_{0}}{\sqrt{\Delta_{0}^{2}+4\Omega_{R}^{2}}}\left(\frac{c}{n\hbar\pi L_{c}^{-1}}-\frac{1}{M_{\text{exc}}}\right)\right]$
The dispersions of the photon, exciton, lower, and upper polaritons is visualised below.
Figure 19: Energy dispersions of the photon, exciton, lower polaritons, and upper polaritons as a function of the in-plane momentum in one of our candidate materials - GaAs. The parameters used were $\Omega_{R}=17$ meV, a bare photon of energy 1.6 eV, and a finely adjusted detuning. The energy is measured relative to the top of the valence band.
The effective masses of the lower and upper polaritons can be found using the effective mass approximation, which relates the inverse effective mass to the second derivative of the energy spectrum with respect to momentum. The general formula is given by
$\frac{1}{m_{\text{eff}}}=\frac{\partial^{2}\varepsilon_{\text{LP/UP}}}{\partial P^{2}}$ (67)
The Taylor-expanded expression for the polariton energy near small momentum $P$ is given by Eq. (66). The term proportional to $P^{2}$ contributes to the effective mass calculation. We denote this coefficient by $A_{\text{LP/UP}}$, where
$A_{\text{LP/UP}}=\frac{1}{4}\left[\left(\frac{c}{n\hbar\pi L_{c}^{-1}}+\frac{1}{M_{\text{exc}}}\right)\mp\frac{\Delta_{0}}{\sqrt{\Delta_{0}^{2}+4\Omega_{R}^{2}}}\left(\frac{c}{n\hbar\pi L_{c}^{-1}}-\frac{1}{M_{\text{exc}}}\right)\right]$ (68)
The effective mass is then
$\frac{1}{m_{\text{LP/UP}}}=2A_{\text{LP/UP}}$ (69)
We find the effective masses for the lower and upper polaritons as
$\frac{1}{m_{\text{LP}}}=\frac{1}{2}\left[\left(\frac{c}{n\hbar\pi L_{c}^{-1}}+\frac{1}{M_{\text{exc}}}\right)-\frac{\Delta_{0}}{\sqrt{\Delta_{0}^{2}+4\Omega_{R}^{2}}}\left(\frac{c}{n\hbar\pi L_{c}^{-1}}-\frac{1}{M_{\text{exc}}}\right)\right]$ (70)
$\frac{1}{m_{\text{UP}}}=\frac{1}{2}\left[\left(\frac{c}{n\hbar\pi L_{c}^{-1}}+\frac{1}{M_{\text{exc}}}\right)+\frac{\Delta_{0}}{\sqrt{\Delta_{0}^{2}+4\Omega_{R}^{2}}}\left(\frac{c}{n\hbar\pi L_{c}^{-1}}-\frac{1}{M_{\text{exc}}}\right)\right]$ (71)
We plot the effective masses of the upper and lower polaritons as a function of the detuning and the Rabi splitting. In Fig. 20 below we plot these dependances.
Figure 20: (Color online) The polariton effective masses in GaAs as a function of (a) the detuning and (b) the Rabi splitting $\Omega_{R}$. For (a), we kept $\Omega_{R}=15$ meV as a constant, and for (b) we set the detuning to be $\Delta_{0}$ = -10 meV.
a) Refer to caption b) Refer to caption
It is worth noting that the effective masses of the upper and lower polariton are constant and identical at a detuning of 0. Away from the detuning of 0, the Rabi splitting allows for the masses of the upper and lower polaritons can be adjusted as a function of the Rabi splitting.
Appendix D Detailed Collective Excitation Spectrum for Two-Component Polaritons
The collective excitation spectrum is given by
$\Omega_{j,\mathbf{P}}=\sqrt{\frac{\omega_{1,\mathbf{P}}^{2}+\omega_{2,\mathbf{P}}^{2}+(-1)^{j-1}\sqrt{\beta}}{2}},$ (72)
where the single-branch frequencies are
$\omega_{j,\mathbf{P}}^{2}=\varepsilon_{j,\mathbf{P}}^{2}+2U_{q}n_{j,0}\varepsilon_{j,\mathbf{P}},$ (73)
with $j=1$ for lower polaritons (LP) and $j=2$ for upper polaritons (UP).
The energy spectrum $\varepsilon_{j,\mathbf{P}}$ is defined as,
$\varepsilon_{j,\mathbf{P}}=\frac{\varepsilon_{\text{ph},\mathbf{P}}+\varepsilon_{\text{exc},\mathbf{P}}}{2}\mp\frac{1}{2}\sqrt{\left(\varepsilon_{\text{ph},\mathbf{P}}-\varepsilon_{\text{exc},\mathbf{P}}\right)^{2}+4\hbar^{2}\Omega_{R}^{2}}.$ (74)
The photon dispersion is given by
$\varepsilon_{\text{ph},\mathbf{P}}=\frac{c}{n}\sqrt{\mathbf{P}^{2}+\left(\frac{\pi}{L_{c}}\right)^{2}}.$ (75)
For small $\mathbf{P}$, this simplifies to,
$\varepsilon_{\text{ph},\mathbf{P}}\approx\frac{c}{n}\left(\frac{\pi}{L_{c}}\right)+\frac{\hbar^{2}\mathbf{P}^{2}}{2m_{\text{ph}}},\quad m_{\text{ph}}=\frac{\hbar\pi}{cnL_{c}}.$ (76)
The exciton dispersion is given by
$\varepsilon_{\text{exc},\mathbf{P}}=E_{\text{band}}-E_{\text{binding}}+\frac{\hbar^{2}\mathbf{P}^{2}}{2m_{\text{exc}}}.$ (77)
The interaction strengths are,
$G_{jj}=U_{q}n_{j,0},$ (78)
where $n_{j,0}$ is the total polariton concentration for species $j$ and $U_{q}$ is the interaction constant. The cross-interaction term is given by:
$G_{12}=U_{q}\sqrt{n_{\text{LP},0}n_{\text{UP},0}}.$ (79)
The coupling and interaction terms contribute to $\beta$, defined as,
$\beta=\left(\varepsilon_{1,\mathbf{P}}^{2}+2U_{q}n_{\text{LP},0}\varepsilon_{1,\mathbf{P}}-\varepsilon_{2,\mathbf{P}}^{2}-2U_{q}n_{\text{UP},0}\varepsilon_{2,\mathbf{P}}\right)^{2}+16U_{q}^{2}n_{\text{LP},0}n_{\text{UP},0}\varepsilon_{1,\mathbf{P}}\varepsilon_{2,\mathbf{P}}.$ (80)
Substituting $\omega_{j,\mathbf{P}}^{2}=\varepsilon_{j,\mathbf{P}}^{2}+2U_{q}n_{j,0}\varepsilon_{j,\mathbf{P}}$, the collective excitation spectrum becomes,
$\Omega_{j,\mathbf{P}}=\sqrt{\frac{\varepsilon_{1,\mathbf{P}}^{2}+\varepsilon_{2,\mathbf{P}}^{2}+2U_{q}n_{\text{LP},0}\varepsilon_{1,\mathbf{P}}+2U_{q}n_{\text{UP},0}\varepsilon_{2,\mathbf{P}}+(-1)^{j-1}\sqrt{\beta}}{2}}.$ (81)