### v2

# Theory of two-component superfluidity of microcavity polaritons

Abstract
We develop a microscopic mean-field theory describing the coexistence of Bose–Einstein condensates of upper and lower polaritons (UP/LP) in a semiconductor microcavity. Incorporating interbranch scattering within a modified polariton Hamiltonian, we introduce a phenomenological population–split parameter $\alpha$ that quantifies the relative LP/UP occupations. At zero detuning, the critical temperature becomes independent of $\alpha$, converging to a single value that marks the balanced, resonant regime. Away from resonance, variations in $\alpha$ lead to distinctive and experimentally resolvable changes in both the sound velocity $c_{s}$ and critical temperature $T_{c}$, relative to the single-component (LP-only) condensate limit. The system under study consists of excitons confined in A transition metal dichalcogenide (TMDC) monolayer, particularly WSe2 embedded within a planar optical microcavity of GaAs where they strongly couple to cavity photons. Our analysis focuses on monolayer WSe2 embdedded in a GaAs microcavity. We present results for GaAs/AlGaAs quantum wells embedded in a GaAs microcavity in the Appendix.While mean-field in scope, the framework provides analytic benchmarks and physical insight for future treatments that include dissipation and fluctuations in nonequilibrium polariton superfluids.
I Introduction
Bose–Einstein condensation (BEC) and superfluidity of exciton-polaritons have become key topics in two-dimensional (2D) systems due to the hybrid light–matter nature of polaritons deng2010 ; snokeKeeling2017 ; kavokin2017 , formed via strong coupling between quantum well excitons and cavity photons Moskalenko2000 ; Littlewood2007 . Excitons are bound electron–hole pairs that form in semiconductors through Coulomb attraction. When such excitons couple strongly to the quantized electromagnetic field of an optical microcavity, they hybridize with cavity photons to form new light–matter quasiparticles known as exciton–polaritons weisbuch1992 . In GaAs/AlGaAs structures, the excitons are confined within a quantum well embedded between distributed Bragg reflectors that define the optical cavity Skol1998 . In transition metal dichalcogenide (TMDC) realizations, by contrast, the excitons reside in an atomically thin monolayer placed inside a similar cavity, where the large binding energy enhances the strength of light–matter coupling dufferwiel2015 . Their low effective mass enables high critical temperatures for superfluidity  Carusotto2013 ; Proukakis2017 ; Peng2022 , with robust behavior under disorder and external fields berman2006 . Polaritons support parametric scattering ciuti2003 ; savona2005 and coherence across diverse materials, including GaAs berman2008 , CdTe richard2005 , and TMDCs manzeli2017 ; lerario2017 , where Bose-Einstein condensation kasprzak2006 and room-temperature superfluidity have been observed dufferwiel2015 ; kogar2017 ; Peng2022 ; zhao2021 , alongside promising optoelectronic applications mak2010 and the emergence of topological excitations in spinor condensates rubo2007 . Theoretical models have further explored nonlinear polariton condensation beyond the linear regime eastham2001 . Additionally, experiments Comaron2025 have highlighted how tunable polariton interactions in driven-dissipative systems enhance spatial coherence and enable Berezinskii–Kosterlitz–Thouless-like crossovers, consistent with interaction-driven superfluid behavior.
The study of polariton condensates naturally extends to a broader class of multi-component quantum fluids that exhibit rich collective phenomena. Multi-component quantum condensates (comprising coexisting, interacting macroscopic quantum states) exhibit collective phenomena absent in single-component superfluids. In spinor atomic Bose–Einstein condensates, simultaneous occupation of multiple hyperfine states gives rise to coupled spin-density modes, relative-phase dynamics, and topological defects such as half-quantum vortices Ho1998 ; Stamper2013 . In magnon systems, multi-branch condensation produces hybridized magnon modes and interaction-mediated phase transitions Demokritov2006 ; Aeppli2022 .An instructive parallel can be drawn with superfluid 3He, where distinct pairing symmetries (A and B phases) correspond to multiple order parameters with coupled collective excitations Vollhardt2003 ; Leggett1975 . Although the microscopic mechanisms differ, the analogy underscores how coexisting condensate components can support richer excitation spectra and phase dynamics than their single-component counterparts.
Microcavity polaritons offer a uniquely tunable platform for exploring two-component condensate physics. The upper (UP) and lower (LP) polariton branches arise naturally from strong exciton–photon coupling, with an energy splitting that can be engineered through cavity design. These two branches correspond to the eigenmodes of the coupled exciton–photon system obtained through the Hopfield diagonalization. The lower polariton (LP) has a larger excitonic fraction near positive detuning, giving it a heavier effective mass and stronger interactions, whereas the upper polariton (UP) becomes more photonic at negative detuning and has a lighter mass. The two modes are separated by the vacuum Rabi splitting $\hbar\Omega_{R}$, which quantifies the strength of the light–matter coupling and sets the energy scale for the polariton system Carusotto2013 ; weisbuch1992 . While most experiments observe condensation exclusively in the lower branch due to efficient phonon-assisted relaxation kasprzak2006 , recent experimental work has demonstrated upper-branch condensation Chen2023 , suggesting that controlled dual-branch occupation may be achievable under appropriate conditions. A key control parameter in such systems is the exciton–photon detuning, $\Delta_{0}$, defined as the energy difference between the bare cavity photon mode and the exciton resonance. Detuning can be tuned in situ by adjusting the cavity length, refractive index (via temperature or angle of incidence), or through fabrication of distributed Bragg reflectors of different thickness dufferwiel2015 . Varying $\Delta_{0}$ changes the Hopfield coefficients that determine the excitonic and photonic fractions of the polariton branches, thereby modifying their effective masses and interaction strengths. Near zero detuning, both components are equally mixed, while large positive or negative detuning yields predominantly excitonic or photonic character, respectively. This tunability, along with control over excitation density and bare Rabi splitting, allows polaritons to access regimes ranging from light-mass, weakly interacting fluids to heavy-mass, strongly interacting excitonic gases.
Despite these experimental advances and the theoretical progress in understanding single-branch polariton condensates, no microscopic theory has yet incorporated mutual coherence and interaction between lower and upper polariton (LP and UP) condensates in a unified superfluid framework. The upper branch, despite experimental signatures of occupation Chen2023 , remains largely absent from such treatments. Previous work has focused largely on one-component polariton condensates, particularly in the lower polariton (LP) branch. Polariton condensates can form in two distinct branches—lower polariton (LP) and upper polariton (UP)—arising from Rabi splitting szymanska2004 ; snokecoherent2006 .
Motivated by advances in multi-component superfluidity and tunable quasiparticle engineering arafat2024semi ; Sanvitto2016 ; alyatkin2020 ; Bhattacharya2013 , we develop a theory of superfluidity in a two-component polariton condensate comprising both LP and UP branches, formulated within an equilibrium mean-field (Bogoliubov) framework that explicitly includes detuning dependence and interaction-driven effects. Our analysis is carried out in the long-wavelength limit, where the lower and upper polariton dispersions are expanded to quadratic order in momentum, and the effective masses treated as constants, allowing analytical access to the collective modes and superfluid parameters such as the concentration of the superfluid component and the critical temperature for superfluidity. Our theoretical framework treats the condensate population distribution between LP and UP branches as a phenomenological input, characterized by a parameter $\alpha\in[0,1]$, and calculates the resulting superfluid properties via mean-field theory. The systems we consider in particular are GaAs/AlGaAs quantum wells embedded in a GaAs microcavity, and TMDC (WSe2) monolayers embedded in a GaAs microcavity regan2022 . Our work establishes theoretical benchmarks for an experimentally relevant regime, soon accessible with current advances in polariton microcavities. We provide quantitative predictions for the collective excitation spectrum, sound velocity, and critical temperature for superfluidity that would uniquely distinguish genuine two-component LP–UP superfluidity from single-branch superfluidity, thereby offering clear signatures for future experimental verification.
[…]
We show that this two-component structure yields enhanced and tunable sound velocity and critical temperature of superfluidity compared to the LP-only case. Our results are general, material-independent, and analytically tractable—offering a predictive framework for future studies of nonequilibrium polariton dynamics, Lindblad-driven condensates, and Josephson-like phase control in strongly coupled light–matter systems.
The remainder of this paper is organized as follows. Section II introduces the modified Hamiltonian for upper and lower polaritons, including the systematic reduction of interaction terms. Section III derives the collective excitation spectrum within the Bogoliubov framework. Section IV examines the tunable sound velocity and its dependence on detuning, Rabi splitting, and the population–split parameter $\alpha$. Section V develops the superfluid criterion and derives analytic expressions for the critical temperature for superfluidity and concentration. In this section we also analyze how the critical temperature for superfluidity depends on the detuning, Rabi splitting, and population-split parameter $\alpha$, and demonstrate a retrieval procedure that allows one to infer $\alpha$ from measured values of the critical temperature $T_{c}$ and critical density $n_{c}$. Section VI concludes with a discussion of the main findings and their implications for tunable two-component polariton superfluidity. Appendices A and B present detailed derivations of the interaction reduction and polariton effective masses that are used in our calculations. Appendix C presents all pertinent results of our analysis for GaAs/AlGaAs embedded in GaAs microcavities.
II Effective Hamiltonian of upper and lower polaritons
In this section, we modify the Hamiltonian for our system of excitons and photons to account for both upper and lower polaritons. By analyzing this modified Hamiltonian, we aim to study the critical temperature for Bose-Einstein condensation and the superfluidity of a two-component polariton system, focusing on the role played by scattering between the polariton branches and the collective excitations in the system.
Polaritons are quantum superpositions of excitons and photons. The total Hamiltonian of excitons and microcavity photons is given by $\hat{H}_{\text{total}}~=~\hat{H}_{\text{exc}}+\hat{H}_{\text{ph}}+\hat{H}_{\text{exc-ph}}$, where $\hat{H}_{\text{exc}}$ is the excitonic Hamiltonian, $\hat{H}_{\text{ph}}$ is the photonic Hamiltonian, and $\hat{H}_{\text{exc-ph}}$ is a Hamiltonian of exciton-photon interaction. We can define each of these explicitly.
We begin with the Hamiltonian for 2D excitons in an infinite homogeneous system, given by
$\hat{H}_{\text{exc}}=\sum_{P}\varepsilon_{\text{ex}}(P)\hat{b}_{P}^{\dagger}\hat{b}_{P}+\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}U_{q}\hat{b}_{P+q}^{\dagger}\hat{b}_{P^{\prime}-q}^{\dagger}\hat{b}_{P}\hat{b}_{P^{\prime}}\ ,$ (1)
where $A$ is a normalization area. Here, the first term represents the energy of the excitons, and the second term represents the exciton-exciton interaction. The operator $\hat{b}_{P}$ is the excitonic annihilation operator, and the operator $\hat{b}^{\dagger}_{P}$ is the excitonic creation operator for momentum $P$. The contact potential $U_{q}$ is approximated by Eq. (12) and from experimental benchmarks. The exciton energy dispersion $\varepsilon_{\text{ex}}(P)$ is given by
$\varepsilon_{\text{ex}}(P)=E_{\text{band}}-E_{\text{binding}}+\varepsilon_{0}(P)\ ,$ (2)
The Hamiltonian for noninteracting microcavity photons is given by
$\hat{H}_{\text{ph}}=\sum_{P}\varepsilon_{\text{ph}}(P)\hat{a}_{P}^{\dagger}\hat{a}_{P}\ \ .$ (3)
Here, the operator $\hat{a}_{P}$ is the photonic annihilation operator, and the operator $\hat{a}^{\dagger}_{P}$ is the photonic creation operator for momentum $P$. For photons confined to a microcavity, the energy dispersion $\varepsilon_{\text{ph}}(P)$ is given by
$\varepsilon_{\text{ph}}(P)=\frac{c}{n}\sqrt{P^{2}+\hbar^{2}\pi^{2}L_{c}^{-2}}\ .$ (4)
where $L_{c}$ is the microcavity length. The exciton-photon interaction term is given by
$\hat{H}_{\text{exc-ph}}=\sum_{P}\hbar\Omega_{R}(\hat{b}_{P}^{\dagger}\hat{a}_{P}+\hat{a}_{P}^{\dagger}\hat{b}_{P}),$ (5)
where $\hbar\Omega_{R}$ is the Rabi splitting, representing the coupling strength between photons and excitons.To fully describe the system, we express the Hamiltonian in the basis of upper and lower polariton operators. We substitute the following expressions from Eq. (6) of Ref. hopfield1958 .
$\hat{b}_{P}=X_{P}\hat{l}_{P}-C_{P}\hat{u}_{P},\quad\hat{b}_{P}^{\dagger}=X_{P}\hat{l}_{P}^{\dagger}-C_{P}\hat{u}_{P}^{\dagger}$ (6)
where $\hat{l}_{P}$ is the lower polariton annihilation operator for momentum $P$, $\hat{u}_{P}$ is the upper polariton annihilation operator for momentum $P$, $X_{P}$ is the excitonic fraction of the lower polariton and $C_{P}$ is the photonic fraction of the lower polariton. The photonic operators may be defined in the basis of upper and lower polariton operators as
$\hat{a}_{P}=C_{P}\hat{l}_{P}+X_{P}\hat{u}_{P},\quad\hat{a}_{P}^{\dagger}=C_{P}\hat{l}_{P}^{\dagger}+X_{P}\hat{u}_{P}^{\dagger}.$ (7)
The photon and exciton fractions in the lower polariton branch, denoted as $X_{P}$ and $C_{P}$ respectively, describe the relative contributions of the photonic and excitonic components in the polariton wavefunction. These fractions are given by
$X_{P}=\frac{1}{\sqrt{1+\left(\frac{\hbar\Omega_{R}}{\varepsilon_{\text{LP}}(P)-\varepsilon_{\text{ph}}(P)}\right)^{2}}},$ (8)
$C_{P}=-\frac{\frac{\hbar\Omega_{R}}{\varepsilon_{\text{LP}}(P)-\varepsilon_{\text{ph}}(P)}}{\sqrt{1+\left(\frac{\hbar\Omega_{R}}{\varepsilon_{\text{LP}}(P)-\varepsilon_{\text{ph}}(P)}\right)^{2}}},$ (9)
where $\varepsilon_{\text{LP}}(P)$ is the energy of the lower polariton branch, $\varepsilon_{\text{ph}}(P)$ is the photon dispersion. The polariton energy is given by
$\varepsilon_{\text{LP/UP}}(P)=\frac{\varepsilon_{\text{ph}}(P)+\varepsilon_{\text{exc}}(P)}{2}\mp\frac{1}{2}\sqrt{\left(\varepsilon_{\text{ph}}(P)-\varepsilon_{\text{exc}}(P)\right)^{2}+4|\hbar\Omega_{R}|^{2}}\ ,$ (10)
The linear part of the Hamiltonian, which includes the linear part of $\hat{H}_{\text{exc}}$ and $\hat{H}_{\text{ph}}$, and the interaction between exciton and photon $\hat{H}_{\text{exc-ph}}$ can be diagonalized by applying unitary transformations and has the form
$\hat{H}_{0}=\sum_{P}\varepsilon_{LP}(P)\hat{l}_{P}^{\dagger}\hat{l}_{P}+\sum_{P}\varepsilon_{UP}(P)\hat{u}_{P}^{\dagger}\hat{u}_{P}$ (11)
The next thing we need to consider are interactions. We substitute the expressions for $\hat{b}_{P}$ and $\hat{b}_{P}^{\dagger}$ into the excitonic interaction term,
$\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}U_{q}\hat{b}_{P+q}^{\dagger}\hat{b}_{P^{\prime}-q}^{\dagger}\hat{b}_{P}\hat{b}_{P^{\prime}}$
where, the contact potential $U_{q}$ for the exciton-exciton interaction is approximated to $U_{0}$, and is defined for GaAs/AlGaAs as Carusotto2013
$U_{0}=\frac{6e^{2}a_{2D}}{\epsilon}$ (12)
and $a_{2D}$ is the effective 2D Bohr radius of excitons.
$a_{2D}=\frac{\hbar^{2}\epsilon}{2\mu_{e-h}e^{2}}$ (13)
It is crucial to note that for monolayer WSe2, the exciton–exciton interaction is more accurately described by the Rytova-Keldysh potential Berkelbach2013 , which accounts for reduced screening in atomically thin materials. Rather than deriving $U_{0}$ microscopically, we adopt the experimentally measured value from polariton blueshift measurements Kravtsov2020 ; flatten2016 . Proceeding we get the following
$\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}U_{0}\left(X_{P+q}\hat{l}_{P+q}^{\dagger}-C_{P+q}\hat{u}_{P+q}^{\dagger}\right)\left(X_{P^{\prime}-q}\hat{l}_{P^{\prime}-q}^{\dagger}-C_{P^{\prime}-q}\hat{u}_{P^{\prime}-q}^{\dagger}\right)\left(X_{P}\hat{l}_{P}-C_{P}\hat{u}_{P}\right)\left(X_{P^{\prime}}\hat{l}_{P^{\prime}}-C_{P^{\prime}}\hat{u}_{P^{\prime}}\right)$ (14)
Expanding this product results in terms involving, pure lower polariton interactions, pure upper polariton interactions, cross-interactions between upper and lower polaritons. We start with the expression
$\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}U_{0}\left(X_{P+q}\hat{l}_{P+q}^{\dagger}-C_{P+q}\hat{u}_{P+q}^{\dagger}\right)\left(X_{P^{\prime}-q}\hat{l}_{P^{\prime}-q}^{\dagger}-C_{P^{\prime}-q}\hat{u}_{P^{\prime}-q}^{\dagger}\right)\left(X_{P}\hat{l}_{P}-C_{P}\hat{u}_{P}\right)\left(X_{P^{\prime}}\hat{l}_{P^{\prime}}-C_{P^{\prime}}\hat{u}_{P^{\prime}}\right)$
The subsequent 16 interaction terms can be combined into a single equation, in a more reasonable order, grouping together terms that are alike. We explicitly sort these terms in Appendix A.
Certain terms in the expanded Hamiltonian represent interactions that are not typical for polaritons due to the following reasons. Polaritons in the upper ($\hat{u}$) and lower ($\hat{l}$) branches generally do not transform into one another without a strong coupling mechanism, especially with the large Rabi splitting. The Rabi splitting represents the energy difference between the upper and lower polariton branches, and when this splitting is significant (on the order of tens of meV), transitions between these branches require a comparable amount of energy, making spontaneous interbranch conversions highly improbable Peng2022 ; Carusotto2013 . As a result, terms where a lower and an upper polariton combine to form two of the same type (either both lower or both upper) are not physically consistent and are sorted through in Appendix A. Processes that would convert one lower and one upper polariton into two lower or two upper polaritons are unlikely to occur without the input of additional energy. This makes certain cross-type conversion terms energetically unfavorable, if even possible, so we discard them from consideration. After removing these non-physical terms, the interaction Hamiltonian simplifies to
$\hat{H}_{\text{int}}=\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}U_{0}\Big[$ $g_{\ell\ell}^{PP^{\prime}q}\,\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}+g_{uu}^{PP^{\prime}q}\,\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}$ (15)
$+g_{\ell u}^{PP^{\prime}q}\,\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}+g_{u\ell}^{PP^{\prime}q}\,\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$+\tilde{g}_{\ell u}^{PP^{\prime}q}\,\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}+\tilde{g}_{u\ell}^{PP^{\prime}q}\,\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}\Big].$
With the Hopfield factor coefficients defined as
$g_{\ell\ell}^{PP^{\prime}q}$ $=X_{P+q}X_{P^{\prime}-q}X_{P}X_{P^{\prime}},$ (16)
$g_{uu}^{PP^{\prime}q}$ $=C_{P+q}C_{P^{\prime}-q}C_{P}C_{P^{\prime}},$ (17)
$g_{\ell u}^{PP^{\prime}q}$ $=X_{P+q}C_{P^{\prime}-q}X_{P}C_{P^{\prime}},$ (18)
$g_{u\ell}^{PP^{\prime}q}$ $=C_{P+q}X_{P^{\prime}-q}C_{P}X_{P^{\prime}},$ (19)
$\tilde{g}_{\ell u}^{PP^{\prime}q}$ $=X_{P+q}C_{P^{\prime}-q}C_{P}X_{P^{\prime}},$ (20)
$\tilde{g}_{u\ell}^{PP^{\prime}q}$ $=C_{P+q}X_{P^{\prime}-q}X_{P}C_{P^{\prime}}.$ (21)
To reiterate, the total Hamiltonian in the upper and lower polariton basis then becomes
$\hat{H}_{\text{total}}$ $=\hat{H}_{0}+\hat{H}_{\text{int}}$ (22)
$=\sum_{P}\varepsilon_{\text{LP}}(P)\,\hat{l}_{P}^{\dagger}\hat{l}_{P}\;+\;\sum_{P}\varepsilon_{\text{UP}}(P)\,\hat{u}_{P}^{\dagger}\hat{u}_{P}$
$\quad+\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}U_{q}\Bigg\{\,g_{\ell\ell}^{PP^{\prime}q}\,\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}$
$\qquad\qquad\qquad+\,g_{uu}^{PP^{\prime}q}\,\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}$
$\qquad\qquad\qquad+\,g_{\ell u}^{PP^{\prime}q}\,\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}$
$\qquad\qquad\qquad+\,g_{u\ell}^{PP^{\prime}q}\,\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$\qquad\qquad\qquad+\,\tilde{g}_{\ell u}^{PP^{\prime}q}\,\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$\qquad\qquad\qquad+\,\tilde{g}_{u\ell}^{PP^{\prime}q}\,\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}\Bigg\}.$
We emphasize that the interaction coefficients $g_{\ell\ell},g_{uu},g_{\ell u},g_{u\ell},\tilde{g}_{\ell u},\tilde{g}_{u\ell}$ defined above are in general momentum-dependent through the Hopfield fractions $X_{P}$ and $C_{P}$. Retaining the full momentum dependence ensures that inter- and intra-branch scattering are treated rigorously across all detunings and Rabi couplings. However, such generality comes at the cost of analytic tractability: closed-form expressions for the Bogoliubov spectrum, sound velocity, and critical temperature become intractable if one carries the momentum dependence throughout. For this reason, we adopt a controlled approximation appropriate for condensates forming near $P=0$. We thus evaluate the Hopfield coefficients at zero in-plane momentum, $X_{P}\to X_{0}$ and $C_{P}\to C_{0}$, and use these values to define effective interaction strengths. This approximation is physically justified because Bose-Einstein condensation occurs predominantly at zero momentum, where the single-particle energies are minimized and superfluid properties such as sound velocity and critical temperature are governed by low-energy, long-wavelength collective modes rather than high-momentum excitations. Additionally, at small $P$, the Hopfield coefficients vary slowly, so $X_{P}\approx X_{0}$ and $C_{P}\approx C_{0}$ remain accurate.
III Spectrum of collective excitations
In analyzing the excitation spectrum of the coexisting LP-UP condensate, we may begin in the long-wavelength limit ($P\to 0$) where the dispersions can be linearized and the effective polariton masses treated as constants. This is the regime directly relevant for superfluid properties such as the sound velocity and the critical temperature. Both quantities are governed by the low-energy, phonon-like Bogoliubov modes rather than by high-momentum single-particle excitations. Moreover, at long wavelengths the momentum dependence of the Hopfield coefficients becomes negligible, allowing the interaction Hamiltonian to be expressed in terms of constant effective couplings AGD2012 . This approximation preserves the essential collective physics while avoiding artifacts introduced by momentum-dependent mass renormalization, and is standard practice when applying the Landau criterion or deriving analytic estimates for $c_{s}$ and $T_{c}$.
We define $\Delta_{0}=\frac{c}{n}\hbar\pi L_{c}^{-1}-\left(E_{\text{band}}-E_{\text{binding}}\right)$ as the detuning. At $P=0$ the Hopfield coefficients become momentum independent,
$X_{P}\to X_{0},\qquad C_{P}\to C_{0},\quad X_{0}^{2}=\tfrac{1}{2}\!\left(1+\frac{\Delta_{0}}{\sqrt{\Delta_{0}^{2}+4(\hbar\Omega_{R})^{2}}}\right),\ C_{0}^{2}=1-X_{0}^{2},$ (23)
Crucially, the Hopfield coefficients retain their full detuning and Rabi-splitting dependence. This allows us to explore off-resonance regimes while maintaining analytic tractability. The resulting expressions for collective modes, sound velocity, and critical temperature are exact within the long-wavelength limit and provide physically faithful benchmarks for comparison with experiments and future numerical treatments that incorporate full momentum dependence. The Hopfield coefficients at $P=0$ can be varied widely by varying the detuning and the Rabi splitting as seen below in Figure 1.
Figure 1: The Hopfield coefficients as a function of the detuning $\Delta_{0}$ defined above for WSe2. We employed the definitions in Eq. (23). We kept the Rabi splitting constant at $\Omega_{R}=$ 8 meV.
For the contact interaction we take $U_{q}\to U_{0}$ using Eq. (12) for GaAs/AlGaAs quantum wells, and the experimentally measured value from polariton blueshift measurements for monolayers of WSe2  Kravtsov2020 ; flatten2016 . We then define the constant effective couplings at $P=0$
$g_{\ell\ell}\equiv U_{0}|X_{0}|^{4},\qquad g_{uu}\equiv U_{0}|C_{0}|^{4},\qquad g_{\ell u}\equiv U_{0}|X_{0}|^{2}|C_{0}|^{2},$ (24)
the interaction Hamiltonian in the polariton basis reduces to
$\hat{H}_{\text{int}}^{(P=0)}=\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}\Big[$ $g_{\ell\ell}\,\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}+g_{uu}\,\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}$ (25)
$+g_{\ell u}\,\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}+g_{\ell u}\,\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$+g_{\ell u}\,\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}+g_{\ell u}\,\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}\Big].$
The non-interacting contribution to the total Hamiltonian is given by Eq. (11), where the single-particle dispersions $\varepsilon_{\mathrm{LP/UP}}$ are defined in Appendix B, Eq. (65). Collecting all terms, we obtain
$\boxed{\hat{H}_{\text{total}}^{(P=0)}=\hat{H}_{0}+\hat{H}_{\text{int}}^{(P=0)}},$ (26)
we obtain the long-wavelength effective Hamiltonian we will use to derive the collective excitation spectrum.
At $T=0$ K we assume that both lower and upper polaritons macroscopically occupy the zero-momentum state, forming a coexisting condensate. Expanding the total Hamiltonian in fluctuations around the condensates and applying the Bogoliubov approximation, we obtain the quadratic Hamiltonian governing the collective excitations. Diagonalization yields two excitation branches with dispersion
$\Omega_{j,P}=\sqrt{\frac{\omega_{\text{LP},P}^{2}+\omega_{\text{UP},P}^{2}+(-1)^{j-1}\sqrt{\left(\omega_{\text{LP},P}^{2}-\omega_{\text{UP},P}^{2}\right)^{2}+\left(4G_{\text{LP-UP}}\right)^{2}\varepsilon_{\text{LP}}(P)\varepsilon_{\text{UP}}(P)}}{2}}\ ,$ (27)
where $j=1,2$ and
$\omega_{\text{LP},P}^{2}$ $=\varepsilon_{\text{LP}}(P)^{2}+2G_{\text{LP}}\,\varepsilon_{\text{LP}}(P),$ (28)
$\omega_{\text{UP},P}^{2}$ $=\varepsilon_{\text{UP}}(P)^{2}+2G_{\text{UP}}\,\varepsilon_{\text{UP}}(P),$ (29)
with interaction parameters
$G_{\text{LP}}$ $=g_{\ell\ell}\,n_{\text{LP},0},\quad g_{\ell\ell}=U_{0}|X_{0}|^{4},$ (30)
$G_{\text{UP}}$ $=g_{uu}\,n_{\text{UP},0},\quad g_{uu}=U_{0}|C_{0}|^{4},$ (31)
$G_{\text{LP-UP}}$ $=g_{\ell u}\,\sqrt{n_{\text{LP},0}n_{\text{UP},0}},\quad g_{\ell u}=U_{0}|X_{0}|^{2}|C_{0}|^{2}.$ (32)
The two-mode Bogoliubov spectrum for coupled condensates has been studied extensively in atomic systems Pindzola , where hybridization between branches arises from interspecies interactions. Here $n_{\mathrm{LP},0}$ and $n_{\mathrm{UP},0}$ denote the polariton densities in the lower and upper polariton branches, $U_{0}$ is the contact exciton-exciton interaction strength from Eq. (12), and $X_{0}$, $C_{0}$ are the Hopfield coefficients evaluated at $P=0$. We plot the resulting collective excitation spectrum in Figure 2 for GaAs quantum wells in a GaAs microcavity. While the Bogoliubov treatment of multi-component condensates is a well-established formalism, its application to the coupled LP–UP polariton system introduces a qualitatively new degree of tunability: both the effective masses and the interaction constants depend explicitly on the Hopfield coefficients. This dependence enables continuous tuning of the condensate’s collective properties through the cavity detuning and Rabi coupling, linking the underlying exciton–photon composition to experimentally observable superfluid characteristics.
Figure 2: The spectrum of collective excitations in GaAs as a function of momentum. The parameters used in the calculation are $\Omega_{R}=17$ meV, a bare photon of energy 1.6 eV, and a detuning $\Delta_{0}=0$ meV. The collective excitation spectrum is defined in Eq. (27).
Eq. (27) describes the collective excitation spectrum in the system with coupled lower- and upper-polariton condensates within the mean-field framework. The interaction coefficients satisfy $G_{\mathrm{LP\!-\!UP}}^{2}=G_{\mathrm{LP}}G_{\mathrm{UP}}$, and the $2\times 2$ interaction matrix becomes degenerate, so that the interaction matrix effectively reduces to a single dominant term. Physically, this condition implies that the two branches predominantly share a common total-density response, while their relative-density channel becomes weakly constrained. Within this simplified equilibrium treatment, the coupled condensates thus support one clearly defined linear (sound-like) excitation associated with the broken global $\mathrm{U}(1)$ symmetry and a second, low-energy mode that is only weakly dispersive. As shown in Figure 2, the $j=1$ branch corresponds to the collective sound-like mode with velocity $c_{1}$ from Eq. (35), while the $j=2$ branch may be interpreted as a relative-phase–like fluctuation between the LP and UP components. We emphasize that identifying this second mode with a distinct symmetry-breaking channel would require a full time-dependent analysis of the coupled order parameters. Such a treatment goes beyond the present static mean-field framework, which only captures equilibrium density responses. In this limit, the second branch should therefore be regarded as a soft collective excitation originating from interbranch coupling, rather than as a rigorously established Goldstone or Leggett mode Wouters2007 ; Leggett2001 . In the long-wavelength limit $P\to 0$, the excitation spectrum becomes linear, $\Omega_{j,P}\approx c_{j}|P|$, with sound velocities
$c_{j}=\sqrt{\frac{G_{\text{LP}}}{2m_{\text{LP}}}+\frac{G_{\text{UP}}}{2m_{\text{UP}}}+(-1)^{j-1}\sqrt{\left(\frac{G_{\text{LP}}}{2m_{\text{LP}}}-\frac{G_{\text{UP}}}{2m_{\text{UP}}}\right)^{2}+\frac{G_{\text{LP-UP}}^{2}}{m_{\text{LP}}m_{\text{UP}}}}}\ ,$ (33)
where $m_{\text{LP}}$ and $m_{\text{UP}}$ are the effective masses of the lower and upper polaritons, respectively. Eq. (33) shows how the Hopfield coefficients enter the collective spectrum through the renormalized couplings $G_{\text{LP}},G_{\text{UP}},$ and $G_{\text{LP-UP}}$. This formulation preserves the full detuning and Rabi splitting dependence of the excitonic and photonic fractions.
The parameterization of the condensate populations requires some care. The Hopfield coefficients $|X|^{2}$ and $|C|^{2}$ characterize the exciton–photon composition of an individual polariton mode, but they do not determine how the total polariton density is distributed between the lower (LP) and upper (UP) polariton branches. In realistic microcavity systems, this population balance is set by kinetic and reservoir processes—such as pumping, relaxation, and decay—or can be directly inferred from experiment. To capture this generally, we introduce a phenomologically motivated dimensionless population-split parameter $\alpha\in[0,1]$, which specifies the fraction of the total polariton density $n_{0}$ residing in the LP and UP branches as follows,
$n_{\mathrm{LP},0}=\alpha n_{0},\qquad n_{\mathrm{UP},0}=(1-\alpha)n_{0}.$ (34)
Once $\alpha$ (or equivalently $n_{\mathrm{LP},0}$ and $n_{\mathrm{UP},0}$) is specified, the Bogoliubov analysis proceeds self-consistently to yield the excitation spectrum, sound velocity, and critical temperature.
Our Bogoliubov analysis remains valid for any externally supplied $(n_{\text{LP},0},n_{\text{UP},0})$ from experiment or kinetic modeling. The subsequent sections explore how $c_{s}$ and $T_{c}$ depend on $\alpha$, providing testable signatures of two-component condensation.
IV Tunable sound velocity
Evaluating the expression for the sound-like branch ($j=1$ branch), and expanding, gives the effective sound velocity
$\boxed{c_{s}=\sqrt{U_{0}\!\left(\frac{|X_{0}|^{4}n_{\mathrm{LP},0}}{m_{\mathrm{LP}}}+\frac{|C_{0}|^{4}n_{\mathrm{UP},0}}{m_{\mathrm{UP}}}\right)}}.$ (35)
This form makes explicit how the sound velocity depends on the Hopfield coefficients, the effective masses of the polariton branches, and the population imbalance between them. We begin by visualising how the sound velocity as a function of detuning for various values of the population split parameter. In this work, the single-branch limits ($\alpha$ = 1 for LP-only and $\alpha$= 0 for UP-only) are modeled as realistic polariton condensates confined to one branch, retaining the Hopfield-dependent effective masses and interaction constants.
Figure 3: Comparison of sound velocity $c_{s}$ in a two-component polariton condensate: (a) dependence on detuning $\Delta_{0}$ at fixed $\Omega_{R}=8$ meV, and (b) dependence on Rabi splitting $\Omega_{R}$ at fixed $\Delta_{0}=5$ meV. In both cases the polariton density is $n=1\times 10^{11}~\text{cm}^{-2}$.
At zero detuning ($\Delta_{0}=0$), the lower and upper polariton fractions are equal ($|X_{0}|^{2}=|C_{0}|^{2}=1/2$), so their contributions to Eq. (35) balance, and the sound velocity converges to a single value for all population splits $\alpha$. This is because at zero detuning, the lower polariton and upper polariton effective masses, $m_{\text{LP}}$ and $m_{\text{UP}}$, stated analytically in Eq. (69)-(70), are equal. Away from resonance ($\Delta_{0}\neq 0$), the Hopfield coefficients $|X_{0}|^{2}$ and $|C_{0}|^{2}$ become asymmetric, and the relative condensate populations shift the overall mean-field energy. As shown in Figure 3(a), this causes the sound velocity to peak at different detunings depending on $\alpha$. For instance, when the condensate is predominantly lower polaritonic ($\alpha=0.75$), the stronger excitonic character ($|X_{0}|^{4}$ term) enhances the effective interaction, leading to a higher $c_{s}$ that peaks near $\Delta_{0}\!\approx\!5~\mathrm{meV}$. Increasing the bare Rabi splitting leads to higher sound velocities, reflected in larger $c_{s}$ at fixed $\Omega_{R}$ for greater $\alpha$. The figure illustrates this trend for discrete coupling strengths and is not intended to imply a continuous variation of $\Omega_{R}$. In Figure 3(b), a slightly positive detuning was chosen, so condensates with a greater lower-polariton fraction exhibit higher sound velocities, reflecting their stronger excitonic contribution to the interaction energy. We would also like to probe the sound velocity as a function of concentration for various total fixed polariton densities $n_{0}$ and $\alpha$ values in Figure (4).
Figure 4: Sound velocity $c_{s}$ in a two-component polariton condensate as (a) a function of the total polariton density $n_{0}$ for different values of the condensate population–split parameter $\alpha$, (b) and as a function of the condensate population–split parameter $\alpha$ for various fixed total polariton densities. The Rabi splitting is fixed at $\Omega_{R}=8~\text{meV}$ and the detuning at $\Delta_{0}=5~\text{meV}$.
The dependence of sound velocity on polariton density is shown in Figure (4)(a), where $c_{s}$ is plotted against $n_{0}$ on logarithmic scales for several values of the population-split parameter $\alpha$. The linear behavior in this log-log representation confirms the $c_{s}\propto\sqrt{n_{0}}$ scaling predicted by Eq. (35). At fixed density, condensates with higher LP fractions ($\alpha\to 1$) exhibit larger sound velocities due to the $|X_{0}|^{4}$ weighting of the LP self-interaction term, which is maximal at positive detunings where the LP branch is predominantly excitonic. Figure 4(b) shows the complementary view of the sound velocity $c_{s}$ versus $\alpha$ at fixed total densities $n_{0}$. At the chosen positive detuning ($\Delta_{0}=5~\mathrm{meV}$), the sound velocity grows monotonically with $\alpha$, interpolating smoothly between the pure UP limit and the pure LP limit for $n_{0}=1\times 10^{11}~\mathrm{cm}^{-2}$. This monotonic $\alpha$-dependence is characteristic of positive detunings, where the lower polariton has a larger excitonic component. For negative detunings, the situation is reversed: the upper polariton becomes more excitonic, leading to a corresponding decrease in $c_{s}$ with increasing $\alpha$.
V Superfluidity
In this section, we determine the criterion for which a weakly interacting Bose gas of two components (upper and lower polaritons) can form a superfluid. We exploit the results from the preceding Section IV. Since at low momenta the energy spectrum of quasiparticles of a weakly interacting gas of polaritons is soundlike, this system satisfies the Landau criterion for superfluidity LifPit . The critical velocity for superfluidity is given by $v_{c}=\text{min}(c_{1},c_{2})$ since the quasiparticles are created at velocities exceeding that of sound for the lowest mode of the quasiparticle dispersion.
The density $\rho_{s}(T)$ of the superfluid is defined as $\rho_{s}(T)=\rho-\rho_{n}(T)$, where $\rho=m_{\text{UP}}n_{\text{UP},0}+m_{\text{LP}}n_{\text{LP},0}$ is the total 2D concentration of the system. Also, $\rho_{n}(T)$ is the density of the normal component. We define the normal component $\rho_{n}(T)$ in the usual way berman2022nanomaterials . We assume that the polariton system moves with velocity u, which means that the superfluid component has velocity u. At finite temperatures $T$, dissipating quasiparticles will emerge in this system. Since their density is small at low temperatures, one may assume that the gas of quasiparticles is an ideal Bose gas. We obtain the density of the superfluid component in the system following the procedure described in Ref. phosphoreneOGR . To calculate the superfluid component density, we begin by defining the mass current J for a Bose gas for quasiparticles in the frame of reference where the superfluid component is at rest by
$J=\int\frac{d^{2}P}{(2\pi\hbar)^{2}}P\left[f[\Omega_{1}(P)-Pu]+f[\Omega_{2}(P)-P\cdot]\right]\ ,$ (36)
where $f[\Omega_{1}(P)]=\{\text{exp}[\Omega_{1}(P)/(k_{B}T)]-1\}^{-1}$ and $f[\Omega_{2}(P)]=\{\text{exp}[\Omega_{2}(P)/(k_{B}T)]-1\}^{-1}$ are the Bose-Einstein distribution functions for the quasiparticles with the angle dependent dispersions $\Omega_{1}(P)$ and $\Omega_{2}(P)$ respectively. Expanding the expression under the integral in terms of $Pu/k_{B}T$ and restricting ourselves to the first-order term, we obtain
$J=-\frac{u}{2}\int\frac{d^{2}P}{(2\pi\hbar)^{2}}P^{2}\left(\frac{\partial f[\Omega_{1}(P)]}{\partial\varepsilon_{1}(P)}+\frac{\partial f[\Omega_{2}(P)]}{\partial\Omega_{2}(P)}\right)\ .$ (37)
Using Eqs. (36) and (37), along with the definition $J=\rho_{n}u$, we define
$\rho_{n}(T)=-\frac{1}{2}\int\frac{d^{2}P}{(2\pi\hbar)^{2}}P^{2}\left(\frac{\partial f[\Omega_{1}(P)]}{\partial\Omega_{1}(P)}+\frac{\partial f[\Omega_{2}(P)]}{\partial\Omega_{2}(P)}\right)\ .$ (38)
Exploiting the methods described in Ref. HighTempSuperOleg for the low temperature case, we obtain the mean field critical temperature
$T_{c}=\left[\frac{2\pi\hbar^{2}\rho c_{s}^{4}}{3\zeta(3)k_{B}^{3}}\right]^{1/3}\ ,$ (39)
where $T_{c}$ is the critical temperature for superfluidity, $\hbar$ is the reduced Planck’s constant, $\rho$ is the total 2D density of the system, accounting for contributions from both polariton branches, $c_{s}$ is the sound velocity defined for the two-component condensate in Eq. (35) and for the one-component condensate are the limits $\alpha=1$ and $\alpha=0$. Additionally, $\zeta(3)$ is the Riemann zeta function evaluated for a value of $3$, and $k_{B}$ is the Boltzmann constant. For the two component condensate of upper and lower polaritons, this is explicitly written in terms of the Hopfield coefficients and effective masses of the upper and lower polartions as follows
$T_{c}=\left[\frac{2\pi\hbar^{2}\rho}{3\zeta(3)k_{B}^{3}}\right]^{1/3}\left(U_{0}{\left(\frac{|X_{0}|^{4}\,n_{\mathrm{LP},0}}{m_{\mathrm{LP}}}\;+\;\frac{|C_{0}|^{4}\,n_{\mathrm{UP},0}}{m_{\mathrm{UP}}}\right)}\right)^{2/3}.$ (40)
We plot the dependance of the critical temperature from Eq. (41) as a function of the detuning, and Rabi splitting.
Figure 5: Critical temperature $T_{c}$ in a two-component polariton condensate in monolayer WSe2. (a) Dependence on detuning $\Delta_{0}$ at fixed Rabi splitting $\Omega_{R}=8$ meV. (b) Dependence on Rabi splitting $\Omega_{R}$ at fixed detuning $\Delta_{0}=5$ meV. In both panels, the total polariton density is held constant at $n_{0}=1\times 10^{11}~\text{cm}^{-2}$, and results are shown for multiple values of the population–split parameter $\alpha$. (c) Contour map of $T_{c}$ as a function of $\Delta_{0}$ and $\Omega_{R}$ at fixed density $n_{0}=1\times 10^{11}~\text{cm}^{-2}$ and population split $\alpha=0.5$ and (d) $\alpha=0.75$. This panel highlights how cavity detuning and light–matter coupling together shape the critical temperature landscape.
At zero detuning ($\Delta_{0}=0$), the lower and upper polariton fractions are equal ($|X_{0}|^{2}=|C_{0}|^{2}=1/2$), so their contributions to Eq. (41) balance, and the critical temperature converges to a single value for all population splits $\alpha$. Away from resonance ($\Delta_{0}\neq 0$), the Hopfield coefficients $|X_{0}|^{2}$ and $|C_{0}|^{2}$ become asymmetric, and the relative condensate populations shift both the sound velocity entering Eq. (41) and the 2D total concentration of the system $\rho$. As shown in Figure 5(a), this causes $T_{c}$ to peak at different detunings depending on $\alpha$. For instance, when the system is predominantly lower polaritonic ($\alpha=0.75$), the enhanced excitonic character increases the interaction-driven sound velocity, which enters as $c_{s}^{4/3}$ in the mean-field expression for $T_{c}$, leading to a higher critical temperature that peaks near $\Delta_{0}\!\approx\!10~\mathrm{meV}$. Increasing the Rabi splitting reduces $T_{c}$ initially for all $\alpha$, before levelling out, with the effect most pronounced for larger values of $\alpha$ where the excitonic contribution to the interaction energy is strongest. In Figure 5, a slightly positive detuning was chosen, so condensates with a greater lower-polariton fraction exhibit higher critical temperatures, reflecting the combined effect of stronger interaction and larger effective mass in determining the superfluid stability.
In addition to the critical temperature, it is convenient to define a critical polariton concentration $n_{c}$ at a given system temperature $T$. This quantity characterizes the minimal density required for superfluidity to emerge at temperature $T$ within our mean-field framework. Starting from the expression for the critical temperature,
$T_{c}=\left[\frac{2\pi\hbar^{2}\rho}{3\zeta(3)k_{B}^{3}}\right]^{1/3}\left[U_{0}\left(\frac{|X_{0}|^{4}\,n_{\mathrm{LP},0}}{m_{\mathrm{LP}}}+\frac{|C_{0}|^{4}\,n_{\mathrm{UP},0}}{m_{\mathrm{UP}}}\right)\right]^{2/3},$ (41)
we substitute the definitions $n_{\mathrm{LP},0}=\alpha n_{0}$ and $n_{\mathrm{UP},0}=(1-\alpha)n_{0}$, and introduce the notation
$S(\Delta_{0},\Omega_{R},\alpha)=\frac{|X_{0}|^{4}}{m_{\mathrm{LP}}}\,\alpha+\frac{|C_{0}|^{4}}{m_{\mathrm{UP}}}\,(1-\alpha).$ (42)
Eq. (41) can then be written as
$T_{c}^{3}=\frac{2\pi\hbar^{2}\rho}{3\zeta(3)k_{B}^{3}}\,\big(U_{0}Sn_{0}\big)^{2}.$ (43)
We also note that the density is defined as,
$\rho=m_{\mathrm{UP}}n_{\mathrm{UP},0}+m_{\mathrm{LP}}n_{\mathrm{LP},0}=n_{0}\Big[m_{\mathrm{LP}}\alpha+m_{\mathrm{UP}}(1-\alpha)\Big]\equiv n_{0}M_{\alpha},$ (44)
where $M_{\alpha}$ is a weighted polariton effective mass determined by the population split parameter $\alpha$. Substituting Eq. (44) for $\rho$ gives
$T_{c}^{3}=\frac{2\pi\hbar^{2}}{3\zeta(3)k_{B}^{3}}\,M_{\alpha}\,(U_{0}S)^{2}\,n_{0}^{3}.$ (45)
Defining
$A\equiv\frac{2\pi\hbar^{2}}{3\zeta(3)k_{B}^{3}},$ (46)
we may write
$T_{c}(\Delta_{0},\Omega_{R},n_{0},\alpha)=\big(AM_{\alpha}\big)^{1/3}\,(U_{0}S)^{2/3}\,n_{0}$} (47)
which shows that $T_{c}$ is linear in the total polariton concentration $n_{0}$ once the density-dependence of $\rho$ is taken into account. Finally, inverting Eq. (47) allows us to identify the critical concentration $n_{c}$ at a given temperature $T$ as
$\boxed{\;n_{c}(T;\Delta_{0},\Omega_{R},\alpha)=\frac{T}{\big(AM_{\alpha}\big)^{1/3}\,(U_{0}S)^{2/3}}\;}.$ (48)
Both $S$ and $M_{\alpha}$ encode the $\alpha$-dependence of the problem. In particular $S$ does so through the Hopfield coefficients and effective polariton masses, and $M_{\alpha}$ through the mass-weighted population imbalance. Thus, measurements of either $T_{c}$ or $n_{c}$ provide a direct probe of the relative occupations of the condensates in the lower and upper polariton branches. The critical concentration $n_{c}$ provides an experimentally accessible quantity that complements the critical temperature $T_{c}$. At fixed detuning $\Delta_{0}$, Rabi splitting $\Omega_{R}$, and interaction strength $U_{0}$, our mean-field framework predicts a one-to-one mapping between $n_{c}$ and the condensate population-split parameter $\alpha$.
This formulation suggests a concrete experimental strategy by measuring the critical density at which superfluidity onsets at a known temperature, an experimentalist can read off the corresponding value of $\alpha$ from our theoretical curves plotted below in Figure (6). This procedure yields two pieces of information simultaneously. The value of $\alpha$, quantifying the relative occupations of the condensates in the lower and upper polariton branches. Additionally it gives confirmation of whether the system indeed realizes a genuine two-component condensate. If the extracted $\alpha$ lies away from the single-component limits ($\alpha=0$ or $\alpha=1$), this serves as direct evidence of coexistence of both branches. Thus, the dependence of $n_{c}$ on $\alpha$ transforms the critical concentration into an experimental probe of population imbalance, offering a pathway to distinguish two-component polariton condensation from its single-component counterparts. An entirely analogous procedure can be applied to the critical temperature $T_{c}$. Measuring $T_{c}$ at fixed density allows one to extract the underlying population imbalance in the same manner.
Figure 6: Condensate population–split parameter $\alpha$ in a two-component polariton condensate in WSe2 monolayers embedded in a GaAs microcavity. (a) Dependence of $\alpha$ on the critical concentration $n_{c}$ at fixed system temperature $T=50$ K, Rabi splitting $\Omega_{R}=8$ meV, and detuning $\Delta_{0}=5$ meV. (b) Dependence of $\alpha$ on the critical temperature $T_{c}$ at fixed total polariton density $n_{0}=1\times 10^{11}~\text{cm}^{-2}$, Rabi splitting $\Omega_{R}=8$ meV, and detuning $\Delta_{0}=10$ meV.
VI Discussion and conclusions
We have developed a microscopic mean-field theory of two-component superfluidity in semiconductor microcavities, treating coexisting condensates of lower and upper polaritons as coupled components of a unified superfluid, with their distinct dispersions, effective masses, and interaction strengths emerging from the underlying exciton-photon Hamiltonian. By incorporating the full momentum and detuning dependence of the Hopfield coefficients, effective masses, and interaction strengths, we derived explicit expressions for the collective excitation spectrum, sound velocity $c_{s}$, and critical temperature $T_{c}$ as functions of the cavity detuning $\Delta_{0}$, Rabi splitting $\Omega_{R}$, and our newly introduced, phenomenological population-split parameter $\alpha$. Our central result is that two-component condensation produces experimentally measurable signatures that distinguish it from conventional single-branch superfluidity, that is enhanced sound velocity and critical temperature across most of the parameter space, distinctive detuning-dependent peak structures in $c_{s}(\Delta_{0},\Omega_{R})$, $T_{c}(\Delta_{0},\Omega_{R})$, and $n_{c}(\Delta_{0},\Omega_{R})$ and a systematic dependence on $\alpha$ that enables direct experimental extraction of the population split from measurements at known system parameters.
At zero detuning, where $|X_{0}|^{2}=|C_{0}|^{2}=1/2$, the system is symmetric and $c_{s}$ and $T_{c}$ converge to single values independent of $\alpha$. Away from resonance, the Hopfield-weighted interaction terms $g_{\ell\ell}=U_{0}|X_{0}|^{4}$, $g_{uu}=U_{0}|C_{0}|^{4}$, and $g_{\ell u}=U_{0}|X_{0}|^{2}|C_{0}|^{2}$ become asymmetric, and both $c_{s}$ and $T_{c}$ develop pronounced $\alpha$-dependence. For instance, at positive detunings in monolayer WSe2 ($\Delta_{0}=5~\mathrm{meV}$, $\Omega_{R}=8~\mathrm{meV}$), an LP-dominated condensate ($\alpha=0.75$) reaches a peak critical temperature $T_{c}\approx 140~\mathrm{K}$, compared to $T_{c}\approx 120~\mathrm{K}$ for a balanced condensate ($\alpha=0.5$). These enhancements arise from the stronger excitonic contribution to the interaction energy when the LP branch is predominantly excitonic in nature.
Many examples demonstrate that multi-component condensation with interbranch coherence is not merely hypothetical but has been realized experimentally in systems with comparable energy splittings and interaction-driven coupling. The magnon case is particularly relevant: Bailey et al. (2024) Aeppli2022 report multi-band Bose-Einstein condensation stabilized by four-particle scattering processes, demonstrating that interaction-mediated coupling can sustain coexisting populations in energetically distinct branches. This experiment was performed in thin films of Yttrium Iron Garnet (YIG), where quantized spin waves (magnons) form quantized subbands due to confinement across the film thickness. Under microwave pumping, Bose–Einstein condensation occurring simultaneously in two distinct magnon bands has been observed. One corresponds to the lowest (Kittel-like) mode and the other to an excited quantized mode. The coexistence of these two condensates is stabilized by strong four-magnon scattering processes that exchange pairs of magnons between the bands, effectively coupling the condensates despite their energy splitting. This establishes a genuine multi-band condensate with interaction-mediated coherence between energetically distinct branches. It is directly analogous to our LP–UP system, where interbranch scattering terms in Eq. (15) play the same coupling role, suggesting that two-component condensation stabilized by interaction-driven coupling is experimentally realistic.
Crucially, our framework is intentionally predictive rather than explanatory. We do not derive the microscopic conditions under which simultaneous LP and UP condensation occurs—that question requires a full kinetic theory incorporating reservoir dynamics, phonon scattering, pump mechanisms, and the conditions for phase-locking between branches—and remains an important target for future work. Instead, we establish theoretical benchmarks that answer the question: if both branches are macroscopically occupied, what collective phenomena emerge, and how can they be unambiguously identified?
Figure 6 provide explicit extraction protocols in that by measuring $c_{s}$ or $T_{c}$ at known $n_{0}$ and $\Delta_{0}$ and $\Omega_{R}$, experimentalists can infer $\alpha$ and thereby confirm the presence and degree of two-component condensation. These predictions are general and apply to any strongly coupled exciton-photon system, with numerical parameters adjusted for specific materials (GaAs/AlGaAs quantum wells and monolayer TMDCs embedded in GaAs microcavities).
While most experiments to date observe single-branch condensation due to efficient phonon-assisted relaxation, recent demonstrations of upper-polariton condensation Chen2023 , suggest that controlled dual-branch occupation may be experimentally accessible under appropriate conditions. These developments echo successful realizations of multi-component condensation in related systems: spinor atomic BECs exhibit simultaneous occupation of multiple hyperfine states with interaction-driven phase-locking Stamper2013 ; Ho1998 , magnon condensates display coexisting branches with hybridized collective modes Aeppli2022 . Our results provide the theoretical framework needed to recognize and characterize analogous phenomena when LP and UP branches are simultaneously macroscopically occupied.
Looking forward, several natural extensions of this work merit investigation. A complete understanding of the coupled LP–UP condensate will ultimately require a time-dependent treatment of the relative phase between the two order parameters, which may give rise to Rabi-like oscillations or other dynamical coherence phenomena beyond the scope of the present equilibrium model. First, a driven–dissipative formulation based on the Lindblad master equation would capture the intrinsically nonequilibrium nature of polariton condensates, incorporating pumping, decay, and reservoir dynamics explicitly to provide time-dependent predictions for $\alpha(t)$ and to determine the conditions under which steady-state dual-branch occupation can emerge. Second, a full amplitude–phase decomposition of the coupled order parameters—together with a microscopic analysis of the phase-locking condition—remains an essential next step. Such a study would establish how interaction-induced coherence and fluctuations determine the stability of the relative phase, thereby connecting the phenomenological population-split parameter $\alpha$ to experimentally controllable pump and reservoir parameters. Finally, our framework naturally generalizes to other multi-component light–matter systems, including cavity magnon-polaritons, phonon-polaritons, and multi-mode photonic condensates, where interaction-driven coupling between branches may yield analogous collective phenomena. By formulating a quantitative equilibrium baseline for two-component polariton superfluidity, this work provides both conceptual motivation and measurable benchmarks for future microscopic and experimental investigations.
Acknowledgements
A. Nafis Arafat and Godfrey Gumbs gratefully acknowledge funding from the U.S. National Aeronautics and Space Administration (NASA) via the NASA-Hunter College Center for Advanced Energy Storage for Space under cooperative agreement 80NSSC24M0177. Oleg Berman acknowledges the support from PSC CUNY Award No. 66382-00 54. This research benefited from Physics Frontier Center for Living Systems funded by the National Science Foundation (PHY- 2317138). The authors are grateful to David W. Snoke for valuable discussions.
Appendix A Systematic interaction reduction
We outline in this section how we systematically reduce Hamiltonian interaction terms to obtain the simplified form in Eq. (22). Let us begin with the 16 interaction terms
$\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}U_{q}\Bigg[$ $X_{P+q}X_{P^{\prime}-q}X_{P}X_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}-X_{P+q}X_{P^{\prime}-q}C_{P}X_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$ (49)
$-X_{P+q}X_{P^{\prime}-q}X_{P}C_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}+X_{P+q}X_{P^{\prime}-q}C_{P}C_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}$
$-X_{P+q}C_{P^{\prime}-q}X_{P}X_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}+X_{P+q}C_{P^{\prime}-q}C_{P}X_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$-X_{P+q}C_{P^{\prime}-q}X_{P}C_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}+X_{P+q}C_{P^{\prime}-q}C_{P}C_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}$
$-C_{P+q}X_{P^{\prime}-q}X_{P}X_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}+C_{P+q}X_{P^{\prime}-q}C_{P}X_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$-C_{P+q}X_{P^{\prime}-q}X_{P}C_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}+C_{P+q}X_{P^{\prime}-q}C_{P}C_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}$
$+C_{P+q}C_{P^{\prime}-q}X_{P}X_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}-C_{P+q}C_{P^{\prime}-q}C_{P}X_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$-C_{P+q}C_{P^{\prime}-q}X_{P}C_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}+C_{P+q}C_{P^{\prime}-q}C_{P}C_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}\Bigg].$
There are only two straightforward self-interaction terms for the upper and lower polaritons. These are
$X_{P+q}X_{P^{\prime}-q}X_{P}X_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}$ (50)
$C_{P+q}C_{P^{\prime}-q}C_{P}C_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}$ (51)
To simplify and rearrange the 16 interaction terms, we will utilize commutation relations for the bosonic creation and annihilation operators AGD2012 , as well as symmetry properties of the terms. The commutation relations reflect the nature of bosons, which can occupy the same quantum state without any restriction. The basic commutation relations for bosonic operators are,
$\hat{l}_{P},\hat{l}_{P^{\prime}}^{\dagger}]=\delta_{P,P^{\prime}}$ (52)
$\hat{u}_{P},\hat{u}_{P^{\prime}}^{\dagger}]=\delta_{P,P^{\prime}}$ (53)
$\hat{l}_{P},\hat{l}_{P^{\prime}}]=0,\quad[\hat{u}_{P},\hat{u}_{P^{\prime}}]=0$ (54)
where $\hat{l}_{P}$ and $\hat{u}_{P}$ are the annihilation operators for different polariton states. Additionally, since these operators correspond to different modes, their cross-commutators will also be zero:
$\hat{l}_{P},\hat{u}_{P^{\prime}}]=[\hat{l}_{P},\hat{u}_{P^{\prime}}^{\dagger}]=0$ (55)
Let us set aside the self interaction terms Eq. (50)-(51) for now. We categorize the 14 remaining cross-interaction terms into two main groups based on the operator structures and then apply normal ordering where possible. Terms Involving $\hat{l}^{\dagger}\hat{l}\hat{u}^{\dagger}\hat{u}$ or $\hat{u}^{\dagger}\hat{u}\hat{l}^{\dagger}\hat{l}$ are interesting to us for considering physical scattering processes.
The terms that fit this structure are
$X_{P+q}C_{P^{\prime}-q}X_{P}C_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P^{\prime}}$
$+C_{P+q}X_{P^{\prime}-q}C_{P}X_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P^{\prime}}$
These two terms involve interactions where two polaritons of different types (upper and lower) scatter with respect to each other. The terms that are remaining involve various other combinations of processes. These remaining 12 cross-interaction terms we categorize into two main groups based on the nature of the exchange processes: direct exchange terms and mixed exchange terms. The direct exchange terms represent direct interactions where the upper polariton exchanges with the lower polariton.
$X_{P+q}C_{P^{\prime}-q}C_{P}X_{P^{\prime}}\,\hat{l}_{P+q}^{\dagger}\,\hat{u}_{P^{\prime}-q}^{\dagger}\,\hat{u}_{P}\,\hat{l}_{P^{\prime}}$
$\quad+C_{P+q}X_{P^{\prime}-q}X_{P}C_{P^{\prime}}\,\hat{u}_{P+q}^{\dagger}\,\hat{l}_{P^{\prime}-q}^{\dagger}\,\hat{l}_{P}\,\hat{u}_{P^{\prime}}$ (56)
These two terms describe simple polariton exchange interactions. The following mixed terms involve more complex combinations of operators that we would like to sort through
$X_{P+q}C_{P^{\prime}-q}X_{P}X_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}$
$+X_{P+q}C_{P^{\prime}-q}C_{P}C_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}$
$+C_{P+q}X_{P^{\prime}-q}C_{P}C_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}$
$+C_{P+q}X_{P^{\prime}-q}X_{P}X_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}$
$+X_{P+q}X_{P^{\prime}-q}C_{P}X_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$+X_{P+q}X_{P^{\prime}-q}X_{P}C_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}$
$+C_{P+q}C_{P^{\prime}-q}X_{P}C_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}$
$+C_{P+q}C_{P^{\prime}-q}X_{P}X_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}$
$+X_{P+q}C_{P^{\prime}-q}C_{P}X_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$+C_{P+q}X_{P^{\prime}-q}X_{P}C_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}$
$+X_{P+q}C_{P^{\prime}-q}C_{P}X_{P^{\prime}}\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$+C_{P+q}X_{P^{\prime}-q}X_{P}C_{P^{\prime}}\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}$ (57)
The terms in Eq. (57) involve three creation/annihilation operators of one type and one of the other. The interactions still involve four polaritons, but the arrangement is asymmetric with respect to ratio of upper and lower polaritons, and extremely unlikely to occur given the Rabi splitting.
Appendix B Expanded Polariton Spectrum and Effective Masses
Polaritons, which result from the strong coupling between excitons and photons, have an energy spectrum that reflects the combined properties of both components. To understand the behavior of polaritons, it is essential to compute their energy spectrum $\varepsilon_{\text{LP/UP}}(P)$, where LP and UP refer to the lower and upper polariton branches, respectively. This spectrum depends on both the photon dispersion $\varepsilon_{\text{ph}}(P)$ and the excitonic dispersion $\varepsilon_{\text{exc}}(P)$.
By analyzing the polariton spectrum in momentum space, we can extract the effective masses of the polaritons, which determine how they behave in real space under external potentials or forces. The effective mass $m_{\text{eff}}$ for polaritons can be obtained by taking the second derivative of the energy spectrum $\varepsilon_{\text{LP/UP}}(P)$ with respect to momentum $P$:
$m_{\text{eff}}=\left(\frac{\partial^{2}\varepsilon_{\text{LP/UP}}(P)}{\partial P^{2}}\right)^{-1}$ (58)
Here, $\varepsilon_{\text{LP/UP}}(P)$ is the energy of the lower or upper polariton, and $P$ is the momentum of the polariton. Once the effective mass is known, we can model the polaritons in real space. The polariton energy is given by
$\varepsilon_{\text{LP/UP}}(P)=\frac{\varepsilon_{\text{ph}}(P)+\varepsilon_{\text{exc}}(P)}{2}\mp\frac{1}{2}\sqrt{\left(\varepsilon_{\text{ph}}(P)-\varepsilon_{\text{exc}}(P)\right)^{2}+4|\hbar\Omega_{R}^{2}|}\ ,$ (59)
Where, $\varepsilon_{\text{LP}}(P)$ and $\varepsilon_{\text{UP}}(P)$ are the lower and upper polariton energy dipsersions, respectively, $\varepsilon_{\text{ph}}(P)$ is the photon energy dispersion, $\varepsilon_{\text{exc}}(P)$ is the exciton energy dispersion, $\Omega_{R}$ is the Rabi splitting, describing the coupling strength between the exciton and photon.
For photons confined to a microcavity, the energy dispersion is given by
$\varepsilon_{\text{ph}}(P)=\frac{c}{n}\sqrt{P^{2}+\hbar^{2}\pi^{2}L_{c}^{-2}}\ .$ (60)
Where, $c$ is the speed of light in vacuum, $n$ is the refractive index of the material inside the cavity, $P$ is the in-plane momentum of the photon, $\hbar$ is the reduced Planck’s constant, $L_{c}$ is the length of the cavity in the direction perpendicular to the plane, $\hbar\pi L_{c}^{-1}$ represents the quantized photon momentum due to the confinement.
The exciton energy dispersion is given by
$\varepsilon_{\text{ex}}(P)=E_{\text{band}}-E_{\text{binding}}+\varepsilon_{0}(P)\ .$ (61)
Where, $E_{\text{band}}$ is the band gap energy, $E_{\text{binding}}$ is the exciton binding energy, $\varepsilon_{0}(P)$ is the kinetic energy of the exciton, given by
$\varepsilon_{0}(P)=\frac{P^{2}}{2M_{\text{exc}}}$ (62)
$M_{\text{exc}}$ is the effective mass of the exciton ($M_{\text{exc}}=m_{e}+m_{h}$), and $P$ is the in-plane momentum of the exciton. The photon energy can be approximated to first-order as
$\varepsilon_{\text{ph}}(P)\approx\frac{c}{n}\left(\hbar\pi L_{c}^{-1}+\frac{P^{2}}{2\hbar\pi L_{c}^{-1}}\right)$ (63)
Now we let $\Delta_{0}=\frac{c}{n}\hbar\pi L_{c}^{-1}-\left(E_{\text{band}}-E_{\text{binding}}\right)$.
The expanded square root term is
$\sqrt{\Delta_{0}^{2}+4\Omega_{R}^{2}+2\Delta_{0}\frac{P^{2}}{2}\left(\frac{c}{n\hbar\pi L_{c}^{-1}}-\frac{1}{M_{\text{exc}}}\right)}\approx\sqrt{\Delta_{0}^{2}+4\Omega_{R}^{2}}+\frac{\Delta_{0}}{\sqrt{\Delta_{0}^{2}+4\Omega_{R}^{2}}}\frac{P^{2}}{2}\left(\frac{c}{n\hbar\pi L_{c}^{-1}}-\frac{1}{M_{\text{exc}}}\right)$ (64)
The Taylor-expanded polariton energies are
$\varepsilon_{\text{LP/UP}}(P)\approx$ $\;\frac{1}{2}\left(\frac{c}{n}\hbar\pi L_{c}^{-1}+E_{\text{band}}-E_{\text{binding}}\right)\mp\frac{1}{2}\sqrt{\Delta_{0}^{2}+4\Omega_{R}^{2}}$ (65)
$+\frac{P^{2}}{4}\left[\left(\frac{c}{n\hbar\pi L_{c}^{-1}}+\frac{1}{M_{\text{exc}}}\right)\mp\frac{\Delta_{0}}{\sqrt{\Delta_{0}^{2}+4\Omega_{R}^{2}}}\left(\frac{c}{n\hbar\pi L_{c}^{-1}}-\frac{1}{M_{\text{exc}}}\right)\right]$
The dispersions of the photon, exciton, lower, and upper polaritons is visualised below.
Figure 7: Energy dispersions of the photon, exciton, lower polaritons, and upper polaritons as a function of the in-plane momentum in one of our candidate materials - GaAs. The parameters used were $\Omega_{R}=17$ meV, a bare photon of energy 1.6 eV, and a finely adjusted detuning. The energy is measured relative to the top of the valence band.
The effective masses of the lower and upper polaritons can be found using the effective mass approximation, which relates the inverse effective mass to the second derivative of the energy spectrum with respect to momentum. The general formula is given by
$\frac{1}{m_{\text{eff}}}=\frac{\partial^{2}\varepsilon_{\text{LP/UP}}}{\partial P^{2}}$ (66)
The Taylor-expanded expression for the polariton energy near small momentum $P$ is given by Eq. (65). The term proportional to $P^{2}$ contributes to the effective mass calculation. We denote this coefficient by $A_{\text{LP/UP}}$, where
$A_{\text{LP/UP}}=\frac{1}{4}\left[\left(\frac{c}{n\hbar\pi L_{c}^{-1}}+\frac{1}{M_{\text{exc}}}\right)\mp\frac{\Delta_{0}}{\sqrt{\Delta_{0}^{2}+4\Omega_{R}^{2}}}\left(\frac{c}{n\hbar\pi L_{c}^{-1}}-\frac{1}{M_{\text{exc}}}\right)\right]$ (67)
The effective mass is then
$\frac{1}{m_{\text{LP/UP}}}=2A_{\text{LP/UP}}$ (68)
We find the effective masses for the lower and upper polaritons as
$\frac{1}{m_{\text{LP}}}=\frac{1}{2}\left[\left(\frac{c}{n\hbar\pi L_{c}^{-1}}+\frac{1}{M_{\text{exc}}}\right)-\frac{\Delta_{0}}{\sqrt{\Delta_{0}^{2}+4\Omega_{R}^{2}}}\left(\frac{c}{n\hbar\pi L_{c}^{-1}}-\frac{1}{M_{\text{exc}}}\right)\right]$ (69)
$\frac{1}{m_{\text{UP}}}=\frac{1}{2}\left[\left(\frac{c}{n\hbar\pi L_{c}^{-1}}+\frac{1}{M_{\text{exc}}}\right)+\frac{\Delta_{0}}{\sqrt{\Delta_{0}^{2}+4\Omega_{R}^{2}}}\left(\frac{c}{n\hbar\pi L_{c}^{-1}}-\frac{1}{M_{\text{exc}}}\right)\right]$ (70)
We plot the effective masses of the upper and lower polaritons as a function of the detuning and the Rabi splitting. In Figure 8 below we plot these dependances.
Figure 8: (Color online) The polariton effective masses in GaAs as a function of (a) the detuning and (b) the Rabi splitting $\Omega_{R}$. For (a), we kept $\Omega_{R}=15$ meV as a constant, and for (b) we set the detuning to be $\Delta_{0}$ = -10 meV.
a) Refer to caption b) Refer to caption
It is worth noting that the effective masses of the upper and lower polariton are constant and identical at a detuning of 0. Away from the detuning of 0, the Rabi splitting allows for the masses of the upper and lower polaritons can be adjusted as a function of the Rabi splitting.
Appendix C Results for GaAs/AlGaAs quantum wells in a GaAs microcavity
We plot in this section all analogous results for GaAs/AlGaAs quantum wells in a GaAs microcavity. We begin by presenting plots of the sound velocity as a function of detuning and the Rabi splitting in Figure 9.
Figure 9: Sound velocity $c_{s}$ of a two-component polariton condensate: (a) dependence on detuning $\Delta_{0}$ at fixed $\Omega_{R}=15$ meV, and (b) dependence on Rabi splitting $\Omega_{R}$ at fixed $\Delta_{0}=20$ meV. The total polariton density is $n_{0}=1\times 10^{11}~\text{cm}^{-2}$.
Similarly we plot the sound velocity as a function of the total polariton density in Figure 10 below.
Figure 10: Sound velocity $c_{s}$ as a function of total polariton density $n_{0}$ for different population-split parameters $\alpha$. The Rabi splitting and detuning are fixed at $\Omega_{R}=15$ meV and $\Delta_{0}=10$ meV, respectively.
Next we present the results of the critical temperature $T_{c}$ as a function of the detuning and Rabi splitting, in Figure (11) below.
Figure 11: Critical temperature $T_{c}$ in a two-component polariton condensate. (a) Dependence on detuning $\Delta_{0}$ at fixed Rabi splitting $\Omega_{R}=15$ meV. (b) Dependence on Rabi splitting $\Omega_{R}$ at fixed detuning $\Delta_{0}=5$ meV. In both panels, the total polariton density is held constant at $n_{0}=1\times 10^{11}~\text{cm}^{-2}$, and results are shown for multiple values of the population–split parameter $\alpha$.
Afte that we present the results of the critical temperature $T_{c}$ as a function of the total polariton density in Figure (12) below.
Figure 12: Critical temperature $T_{c}$ in a two-component polariton condensate as a function of the total polariton density $n_{0}$ for multiple values of the condensate population–split parameter $\alpha$. The Rabi splitting is fixed at $\Omega_{R}=15$ meV and the detuning at $\Delta_{0}=5$ meV.
Finally we present the phenomenological inference plots for GaAs/AlGaAs below in Figure 13.
Figure 13: Condensate population–split parameter $\alpha$ in a two-component polariton condensate. (a) Dependence of $\alpha$ on the critical concentration $n_{c}$ at fixed system temperature $T=50$ K, Rabi splitting $\Omega_{R}=15$ meV, and detuning $\Delta_{0}=5$ meV. (b) Dependence of $\alpha$ on the critical temperature $T_{c}$ at fixed total polariton density $n_{0}=1\times 10^{11}~\text{cm}^{-2}$, Rabi splitting $\Omega_{R}=8$ meV, and detuning $\Delta_{0}=10$ meV.